### v2

# Quantitative Contraction Rates for McKean-Vlasov Stochastic Differential Equations with Multiplicative Noise

Abstract
This work focuses on the quantitative contraction rates for McKean-Vlasov stochastic differential equations (SDEs) with multiplicative noise. Under suitable conditions on the coefficients of the SDE, this paper derives explicit quantitative contraction rates for the convergence in Wasserstein distances of McKean-Vlasov SDEs using the coupling method. The contraction results are then used to prove a propagation of chaos uniformly in time, which provides quantitative bounds on convergence rate of interacting particle systems, and establishes exponential ergodicity for McKean-Vlasov SDEs.
Keywords. McKean-Vlasov stochastic differential equations, contraction, exponential ergodicity, coupling, propagation of chaos.
Mathematics Subject Classification. 60J25, 60H10, 60J60.
1 Introduction
We consider the McKean-Vlasov stochastic differential equation (SDE)
$\mathrm{d}X_{t}$ $=b(X_{t},\mu_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}W_{t},$ (1.1)
$\mu_{t}$ $=\mathcal{L}(X_{t}),$
where $W$ is a $d$-dimensional Brownian motion, $b:\mathbb{R}^{d}\times\mathcal{P}_{1}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$, and $\mu_{t}=\mathcal{L}(X_{t})$ is the law of the random variable $X_{t}$. The precise conditions on $b$ and $\sigma$ guaranteeing (1.1) has a unique strong solution will be given in what follows. Equation (1.1) is an SDE whose drift coefficient depends not only on the process itself but also on the probability distribution of the process at time $t$. Note also that the diffusion coefficient of (1.1) is state-dependent. This type of equation naturally arises as the limit of system of interacting “particles.” Consider, for example, a large number of players with symmetric dynamics whose positions depend on the positions of the mean fields of other players. As the number of players tends to infinity, the limit dynamic of each player does not depend on the positions of the others anymore but only on their statistical distributions thanks to the law of large numbers. The resulting limit system is a McKean-Vlasov process described by equation (1.1). The study of McKean-Vlasov SDEs was initiated in McKean, (1966, 1967) and further developed in Funaki, (1984), Sznitman, (1991). Since the introduction of mean-field games independently by Huang et al., (2006) and Lasry and Lions, (2007), McKean-Vlasov SDEs have received growing attention; some of the recent developments in this direction can be found in Ding and Qiao, (2021), Chaudru de Raynal, (2020), Wang, (2018) and the references therein. We also refer to the survey paper Huang et al., (2021) for an extensive review of the recent progresses in the study of McKean-Vlasov SDEs.
The investigation of convergence to equilibrium of Markov processes is of central interest and vital importance. Crucial questions include: When does an invariant measure exist? Is it unique? What is the convergence rate? For classical SDEs, these questions have been extensively investigated in the literature. Many different methods have been developed to approach these questions as well. In particular, the classical Harris Theorem gives geometric ergodicity under a minimization and a Lyapunov drift conditions. We refer to Hasminskiĭ, (1960), Khasminskii, (2012) for diffusion processes and Meyn and Tweedie, 1993a , Meyn and Tweedie, 1993b for general Markov processes; see also Hairer and Mattingly, (2011), Hairer et al., (2011) for some new perspectives of the Harris Theorem. The coupling method is a simple but powerful probabilistic tool for the study of convergence of Markov processes. The classical treatments of the coupling method for diffusion processes can be found in Lindvall and Rogers, (1986), Chen and Li, (1989), Priola and Wang, (2006) among others and was further developed in Shao and Xi, (2013), Bao et al., (2014), Cloez and Hairer, (2015), Nguyen and Yin, (2018) for regime-switching diffusion processes, Liang and Wang, (2020), Majka, (2017) for stochastic differential equation driven by Lèvy processes, and Bao et al., (2014), Wang et al., (2022) for functional stochastic differential equations. The books Chen, (2004) and Lindvall, (2002) contain unified and comprehensive treatment of the coupling method.
Nevertheless, little is known for the convergence of McKean-Vlasov processes partly due to the delicate and subtle distribution dependence in their dynamics. Hammersley et al., (2021) presents a sufficient condition for the existence of an invariant measure for McKean-Vlasov SDEs, but the issues of uniqueness and convergence rate were not addressed in the paper. It should also be mentioned that Wang, (2018) derives exponential ergodicity in Wasserstein-2 distance for general McKean-Vlasov SDEs under certain dissipative conditions. In contrast to Wang, (2018), this work aims to quantify the contraction rate to the invariant measure of (1.1) in Wasserstein-1 and a multiplicative Wasserstein distances using the coupling method under contractivity and dissipativity. However, owing to consideration of the contraction rate we need a slightly stronger condition with the diffusion coefficient not depending on the distribution. This allows us to construct an appropriate coupling. While the coupling method has been well-developed for classical SDEs, care must be taken when dealing with McKean-Vlasov SDEs. As pointed out in Eberle et al., (2019) and Hammersley et al., (2021), solutions to (1.1) with different initial laws follow different dynamics with different drift and diffusion coefficients. Consequently, even if one can find a finite coupling time $T$, it is not clear how one can construct a coupling $(X_{t},Y_{t})$ satisfying $X_{t}=Y_{t}$ for all $t\geq T$.
We adopt the coupling method in Eberle et al., (2019), which deals with classical SDEs and McKean-Vlasov SDEs with additive noise, to derive quantitative Harris-type theorems for McKean-Vlasov processes with multiplicative noise and general drift coefficients. Using a similar coupling method as that of Eberle et al., (2019) we obtain explicit quantitative contraction rates for the McKean-Vlasov SDE given in (1.1). The essence of our approach uses a mixed coupling. Such a method uses a synchronous coupling when the two marginal processes are close to each other and a reflection coupling when they are farther apart. Such an idea was also used in Zimmer, (2017) for infinite-dimensional SDEs with additive noise. Our first result, Theorem 3.3, shows a contraction in the Wasserstein-1 distance under contractivity at infinity along with certain perturbation bounds on the diffusion coefficient $\sigma$. This result is then used to derive a uniformly in time propagation of chaos bound in Proposition 4.2 and further obtain the existence of a unique invariant measure and exponential ergodicity with an additional growth condition on the drift $b$ in Theorem 4.4. We next present a local contraction result using a multiplicative Wasserstein distance that depends on the initial law under a dissipative drift condition in Theorem 3.6. This contraction result also leads to a uniform in time propagation of chaos bound along with the exponential ergodicity for the process (1.1); see Proposition 4.3 and Theorem 4.5. In Theorems 3.3 and 3.6, the contraction rates can be computed from the coefficients $b$ and $\sigma$ explicitly.
While both Theorems 3.3 and 3.6 are derived using the coupling method, it is worth pointing out the difference between the contraction at infinity and the dissipative drift conditions. Roughly, contraction at infinity pertains to the interplay among the net drift of the marginal processes, which tends to push them closer to each other when they diverge far apart. The dissipative condition, however, ensures the marginal process is steered back towards the origin if it strays too far from it.
The paper is organized as follows. Section 2 introduces the coupling for (1.1), which is essential for our analyses throughout the paper. Several technical lemmas are also presented in Section 2. Section 3 presents our main contraction results. These contraction results are used to derive a propagation of chaos result as well as exponential ergodicity for (1.1) in Section 4. Appendix A contains the proofs of several lemmas.
Throughout this paper, we denote by $\langle\cdot,\cdot\rangle$ and $|\cdot|$ the Euclidean inner product and the corresponding norm, respectively. Furthermore, $\|\cdot\|$ is the matrix norm induced by the Euclidean norm. If $(E,\rho)$ is a Polish space, $\mathcal{P}(E)$ is the collection of probability measures on $E$ and $\mathcal{P}_{1}(E)$ is the subset of $\mathcal{P}(E)$ with finite first absolute moment. Furthermore, we write $\mu(|\cdot|):=\int_{\mathbb{R}^{d}}|x|\mu(dx)$ if $\mu\in\mathcal{P}_{1}(\mathbb{R}^{d})$. We define as in Villani, (2009) the Wasserstein distance of two probability measures $\mu,\nu\in\mathcal{P}(E)$ on a Polish metric space $(E,\rho)$ by
$\mathcal{W}_{\rho,p}(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\left[\int_{E\times E}\rho(x,y)^{p}\pi(\mathrm{d}x,\mathrm{d}y)\right]^{1/p},$
where $p\in[0,\infty)$ and $\Pi(\mu,\nu)$ is the collection of coupling measures, i.e., $\pi\in\Pi(\mu,\nu)$ for $\mu,\nu\in\mathcal{P}(E)$ if $\pi\in\mathcal{P}(E\times E)$, $\pi(A\times E)=\mu(A)$ and $\pi(E\times A)=\nu(A)$ for every Borel set $A\subset E$. When $p=1$ we simply write $\mathcal{W}_{\rho,1}(\mu,\nu)=\mathcal{W}_{\rho}(\mu,\nu)$. In fact, the Wasserstein distance is well defined if $\rho$ is only a semi-metric. For the case $(E,\rho)=(\mathbb{R}^{d},|\cdot|)$, we write
$\mathcal{W}_{p}(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\left[\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{p}\pi(\mathrm{d}x,\mathrm{d}y)\right]^{1/p},\text{ for }p\geq 1.$
1.1 Assumptions
We use the following assumptions throughout the paper.
Assumption 1.1 (Generalized Lipschitz condition on $b$).
There exists a continuous and bounded function $\kappa:[0,\infty)\to\mathbb{R}$ and a constant $L_{1}\geq 0$ such that
$\langle x-y,b(x,\mu)-b(y,\nu)\rangle\leq\kappa(|x-y|)|x-y|^{2}+L_{1}\mathcal{W}_{1}(\mu,\nu)|x-y|,\quad\forall x,y\in\mathbb{R}^{d},\ \mu,\nu\in\mathcal{P}_{1}(\mathbb{R}^{d}).$
Assumption 1.2 (Lipschitz condition on $\sigma$).
There exist a constant $L_{2}\geq 0$ such that
$\|\sigma(x)-\sigma(y)\|^{2}$ $\leq L_{2}|x-y|^{2}.$ (1.2)
Assumption 1.3 (Growth condition on $b$).
There exists a constant $L_{3}>0$ such that
$|b(0,\mu)|\leq L_{3}(1+\mu(|\cdot|)).$
Assumption 1.4 (Boundedness of $\sigma$ and $\sigma^{-1}$).
The diffusion matrix $\sigma$ is uniformly bounded and nondegenerate in that
$\|\sigma(x)-\sigma(y)\|\leq M\quad\text{ and }\quad\|\sigma(x)^{-1}\|\leq\Lambda,\qquad\forall x,y\in\mathbb{R}^{d},$ (1.3)
where $\Lambda>0$ is a constant.
Assumption 1.5.
The constants $M$ and $\Lambda$ in Assumption 1.4 satisfy
$D:=2\Lambda^{-1}-M>0.$ (1.4)
Remark 1.6.
We note that for any initial condition $\nu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d})$, (1.1) has a unique strong non-explosive solution under Assumptions 1.1, 1.2, and 1.3; see, for example, Wang, (2018). The boundedness and nondegeneracy condition (1.3) is imposed so that we can derive the exponential contractions in Section 3.
Remark 1.7.
The function $\kappa$ in assumption 1.1 may take negative values. Our first main result will require that $\kappa$ is negative for large inputs.
Finally, since $\kappa$ is continuous, we define
$\kappa_{r}:=\sup\{|\kappa(s)|:s\in[0,r]\},\quad\forall r>0.$ (1.5)
2 Preliminary Results
We construct a coupling similar to that in Eberle et al., (2019) using the reflection coupling of Lindvall and Rogers, (1986). We start by recalling the synchronous and reflection couplings for the classical SDE
$\mathrm{d}X_{t}=b(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t}.$ (2.1)
Given initial values $(x_{0},y_{0})\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ and a $d$-dimensional Brownian motion $B_{t}$, the synchronous coupling of two solutions to (2.1) is a diffusion process $(X_{t},Y_{t})\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ solving
$\mathrm{d}X_{t}$ $=b(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},\ \ \$$X_{0}=x_{0},$
$\mathrm{d}Y_{t}$ $=b(Y_{t})\mathrm{d}t+\sigma(Y_{t})\mathrm{d}B_{t},$ $Y_{0}=y_{0}.$
The reflection coupling of two solutions to (2.1) is a diffusion process $(X_{t},Y_{t})$ with values in $\mathbb{R}^{2d}$ solving
$\mathrm{d}X_{t}$ $=b(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},\ \ \ (X_{0},Y_{0})=(x_{0},y_{0})$
$\mathrm{d}Y_{t}$ $=b(Y_{t})\mathrm{d}t+\sigma(Y_{t})H_{t}\mathrm{d}B_{t},\ \ \ \text{for}\ t<T,\ \quad\text{ and }\ \quad X_{t}=Y_{t},\ \ \ \text{for}\ t\geq T,$
where $T=\inf\{t\geq 0:X_{t}=Y_{t}\}$ is the coupling time, and the matrix $H_{t}$ is defined as $H_{t}=I-2u_{t}u_{t}^{\top}$ with $u_{t}=(\sigma(Y_{t}))^{-1}(X_{t}-Y_{t})/|\sigma(Y_{t})^{-1}(X_{t}-Y_{t})|$ for $t<T$.
Heuristically, the reflection coupling constructs a ”mirror hyperplane” between the two processes $X_{t}$ and $Y_{t}$. The idea is that the Brownian motion should eventually help the process hit the mirror, at which point the two processes coincide. Once the processes coincide, we can run them together by switching to the synchronous coupling. Notice that this idea will not work in the case where the drift $b$ has measure dependence, as we would require both $X_{t}=Y_{t}$ and $\mu_{t}=\nu_{t}$, which will not be guaranteed in the above construction. Nevertheless, we can still use this idea to construct a combined coupling.
We now construct a coupling for (1.1), which is adapted from that in Eberle et al., (2019). Roughly, the idea is to fix some $\delta\in(0,1)$ and use the reflection coupling when $|X_{t}-Y_{t}|\geq\delta$, and the synchronous coupling when $|X_{t}-Y_{t}|\leq\delta/2$; in between we use a mixture of both couplings.
To proceed, we define Lipschitz functions $\mathrm{rc}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to[0,1]$ and $\mathrm{sc}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to[0,1]$ satisfying
$\mathrm{sc}^{2}(x,y)+\mathrm{rc}^{2}(x,y)=1.$ (2.2)
Furthermore, we fix some $\delta\in(0,1)$ and impose $\mathrm{rc}(x,y)=1$ for $|x-y|\geq\delta$ and $\mathrm{rc}(x,y)=0$ for $|x-y|\leq\delta/2$. Note this implies that $\mathrm{sc}(x,y)=0$ for $|x-y|\geq\delta$ and $\mathrm{sc}(x,y)=1$ for $|x-y|\leq\delta/2$. Now, we note that under Assumptions 1.1, 1.2, and 1.3, there exists a unique, non-explosive solution $(X_{t})$ to (1.1) for any initial probability measure $\mu_{0}$ by (Wang, , 2018, Theorem 2.1). Therefore, we can define for a fixed initial probability measure $\mu_{0}$,
$b_{\mu_{0}}(t,x)=b(x,\mu_{t}),$
where $\mu_{t}=\mathcal{L}_{X_{t}}$ is the law of $X_{t}$. Moreover, Assumption 1.1 implies that $b_{\mu_{0}}$ is Lipschitz continuous with respect to $x$. Recall that $\sigma$ is Lipschitz continuous thanks to (1.2). Thus, by (Karatzas and Shreve, , 1991, Chapter 5, Theorem 2.5), for fixed probability measures $\mu_{0}$ and $\nu_{0}$, parameter $\delta>0$, and independent Brownian motions $(B_{t}^{1})$ and $(B_{t}^{2})$, we can construct drift coefficients $b_{\mu_{0}}$ and $b_{\nu_{0}}$ and define the coupling $(U_{t})=(X_{t},Y_{t})$ as the unique, strong, and non-explosive solution to
$\mathrm{d}X_{t}$ $=b_{\mu_{0}}(t,X_{t})\mathrm{d}t+\sigma(X_{t})[\mathrm{sc}(U_{t})\mathrm{d}B_{t}^{1}+\mathrm{rc}(U_{t})\mathrm{d}B_{t}^{2}],$ (2.3)
$\mathrm{d}Y_{t}$ $=b_{\nu_{0}}(t,Y_{t})\mathrm{d}t+\sigma(Y_{t})[\mathrm{sc}(U_{t})\mathrm{d}B_{t}^{1}+\mathrm{rc}(U_{t})H_{t}\mathrm{d}B_{t}^{2}],$
with $(X_{0},Y_{0})=(x_{0},y_{0})$ and $H_{t}=I-2u_{t}u_{t}^{\top},$ where
[…]
Note that the choice of $u_{t}$ when $X_{t}=Y_{t}$ is irrelevant since $\mathrm{rc}(U_{t})=0$ in that case. Now note that $H_{t}H_{t}^{\top}=I$ and so combined with (2.2), using Levy’s characterization of Brownian motion, $(X_{t})$ and $(Y_{t})$ solve
$\mathrm{d}X_{t}$ $=b_{\mu_{0}}(t,X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}B_{t},\ \ \ \ X_{0}=x_{0},$ (2.4)
$\mathrm{d}Y_{t}$ $=b_{\nu_{0}}(t,Y_{t})\mathrm{d}t+\sigma(Y_{t})\mathrm{d}\hat{B}_{t},\ \ \ \ \ \ Y_{0}=y_{0},$ (2.5)
with respect to the Brownian motions
$B_{t}$ $=\int_{0}^{t}\mathrm{rc}(U_{s})\mathrm{d}B_{s}^{1}+\int_{0}^{t}\mathrm{sc}(U_{s})\mathrm{d}B_{s}^{2},\ \ \ \ \ \text{and}$
$\hat{B}_{t}$ $=\int_{0}^{t}\mathrm{rc}(U_{s})H_{s}\mathrm{d}B_{s}^{1}+\int_{0}^{t}\mathrm{sc}(U_{s})\mathrm{d}B_{s}^{2}.$
Assumptions 1.1 and 1.2 guarantee the solutions to (2.4) and (2.5) are pathwise unique. Thus they coincide almost surely with the solution of (1.1) with respect to the Brownian motions $(B_{t})$ and $(\hat{B}_{t})$ above and initial values $x_{0}$ and $y_{0}$, respectively. Hence $(X_{t},Y_{t})$ is a coupling for (1.1).
Consider the coupling $U_{t}=(X_{t},Y_{t})$ as written in (2.3). Now define the process $Z_{t}=X_{t}-Y_{t}$ and the radial process $r_{t}=|Z_{t}|$. Then we have
$\mathrm{d}Z_{t}$ $=(b_{\mu_{0}}(t,X_{t})-b_{\nu_{0}}(t,Y_{t}))\mathrm{d}t+\mathrm{sc}(U_{t})(\sigma(X_{t})-\sigma(Y_{t}))\mathrm{d}B_{t}^{1}+\mathrm{rc}(U_{t})(\sigma(X_{t})-\sigma(Y_{t})H_{t})\mathrm{d}B_{t}^{2}$
$=\beta_{t}\mathrm{d}t+\mathrm{sc}(U_{t})\Delta_{t}\mathrm{d}B_{t}^{1}+\mathrm{rc}(U_{t})\alpha_{t}\mathrm{d}B_{t}^{2},$ (2.6)
where throughout the paper, we denote $\Delta_{t}=\sigma(X_{t})-\sigma(Y_{t})$, $\alpha_{t}=\sigma(X_{t})-\sigma(Y_{t})H_{t}$, and $\beta_{t}=b_{\mu_{0}}(t,X_{t})-b_{\nu_{0}}(t,Y_{t})$ for notational simplicity. In addition, the dynamics for the radial process $r_{t}$ is described by the following lemma whose proof is deferred to Appendix A.
Lemma 2.1.
Suppose Assumptions 1.1 and 1.2, then we have
$\mathrm{d}r_{t}=$ $I_{\{r_{t}>0\}}\bigg[\langle e_{t},\beta_{t}\rangle\mathrm{d}t+\frac{\mathrm{rc}^{2}(U_{t})}{2r_{t}}(\mathrm{tr}(\alpha_{t}\alpha_{t}^{\top})-|\alpha_{t}^{\top}e_{t}|^{2})\mathrm{d}t+\frac{\mathrm{sc}^{2}(U_{t})}{2r_{t}}(\mathrm{tr}(\Delta_{t}\Delta_{t}^{\top})-|\Delta_{t}^{\top}e_{t}|^{2})\mathrm{d}t$
$+\mathrm{sc}(U_{t})\langle e_{t},\Delta_{t}\mathrm{d}B_{t}^{1}\rangle+\mathrm{rc}(U_{t})\langle e_{t},\alpha_{t}\mathrm{d}B_{t}^{2}\rangle\bigg]$ (2.7)
almost surely, where
$e_{t}=\begin{cases}\frac{Z_{t}}{r_{t}},&\text{ if }r_{t}>0,\\[4.30554pt] \frac{b_{\mu_{0}}(t,X_{t})-b_{\nu_{0}}(t,Y_{t})}{|b_{\mu_{0}}(t,X_{t})-b_{\nu_{0}}(t,Y_{t})|},&\text{ if }r_{t}=0\text{ and }b_{\mu_{0}}(t,X_{t})-b_{\nu_{0}}(t,Y_{t})\neq 0,\\[4.30554pt] e,&\text{ otherwise},\end{cases}$
in which $e$ is an arbitrary unit vector.
The following lemma will be used to bound the drift when $r_{t}$ is small. The proof of the lemma, which consists of elementary calculations, is arranged in Appendix A.
Lemma 2.2.
For $x\neq y\in\mathbb{R}^{d}$, denoting z=xyz=x-y, $e=z/|z|$, $u(x,y)=(\sigma(y)^{-1}z)/|\sigma(y)^{-1}z|$, $H(x,y)=I-2u(x,y)u(x,y)^{\top}$, $\Delta(x,y)=\sigma(x)-\sigma(y)$, and $\alpha(x,y)=\sigma(x)-\sigma(y)H(x,y)$. Then
$\mathrm{tr}(\alpha(x,y)\alpha(x,y)^{\top})-|\alpha(x,y)^{\top}e|^{2}=\mathrm{tr}(\Delta(x,y)\Delta(x,y)^{\top})-|\Delta(x,y)^{\top}e|^{2}.$ (2.8)
In addition, under Assumption 1.2, we have
$\big|\mathrm{tr}(\Delta(x,y)\Delta(x,y)^{\top})-|\Delta(x,y)^{\top}e|^{2}\big|\leq 2dL_{2}|x-y|^{2}.$ (2.9)
Using Lemma 2.2 with Assumptions 1.2 and 1.4, we can define a bound $A$ by
$A:=\sup_{x\neq y}\bigg\{\frac{d\|\sigma(x)-\sigma(y)\|^{2}|x-y|^{2}-|(\sigma(x)-\sigma(y))^{\top}(x-y)|^{2}}{|x-y|^{3}}\bigg\}$ (2.10)
where $d$ is the dimension. Apparently, $A\geq 0$. We also claim that $A<\infty$. Indeed, since $\|\sigma(x)-\sigma(y)\|\leq M$ thanks to Assumption 1.4, the expression $\frac{d|\sigma(x)-\sigma(y)|^{2}|x-y|^{2}-|(\sigma(x)-\sigma(y))^{\top}(x-y)|^{2}}{|x-y|^{3}}$ in (2.10) is bounded when |xy||x-y| is large. On the other hand, Lemma 2.2 ensures that it is also bounded when |xy||x-y| approaches 0. Therefore the claim that $A<\infty$ follows.
3 Main Results
This section gives two different contraction rates for two solutions starting with different initial values. The first result assumes that as the two processes move far away from each other, the net drift between them is strong enough to push them back together. In this sense we ensure that the function $\kappa$ is negative enough as $r$ gets large. The second result assumes that the drift pushes radially inward when the process is far away from the origin. Here we ensure that the coefficient $b$ is negative enough when $X_{t}$ is large.
3.1 Contraction Result Under Contractivity at Infinity
Our first result needs the following assumption.
Assumption 3.1.
The function $\kappa(r)$ in Assumption 1.1 satisfies
$\limsup_{r\to\infty}\kappa(r)<0.$ (3.1)
In order to construct the metric $\rho$ to be used in Theorem 3.3, we define the constants $R_{1}$ and $R_{2}$ to be
$R_{1}$ $:=1\vee\inf\{R\geq 0:\kappa(r)r<-A,\ \forall r\geq R\},$
$R_{2}$ $:=\inf\left\{R>R_{1}:\kappa(r)\leq-\left(\frac{D^{2}}{R(R-R_{1})}+\frac{A}{R}\right),\ \ \forall r\geq R\right\},$
where $A$ is the nonnegative constant defined in (2.10) and $D$ is defined in (1.4). Note that $R_{1},R_{2}<\infty$ by (3.1). Next we define
$\varphi(r)=\exp\left(-\frac{2}{D^{2}}\int_{0}^{r}(u\kappa^{*}(u)+A)\mathrm{d}u\right),\ \text{ and }\ \Phi(r)=\int_{0}^{r}\varphi(s)\mathrm{d}s,\quad\forall r\geq 0,$
where $\kappa^{*}$ is defined by $\kappa^{*}(r)=\kappa(r)$ when $r\kappa(r)\geq-A$ and $\kappa^{*}(r)=-A/r$ when $r\kappa(r)<-A$. Clearly we have $\kappa^{*}(r)\geq\kappa(r)$ and $r\kappa^{*}(r)+A\geq 0$ for all $r\geq 0$. Hence $\varphi(r)\leq 1$ for $r\geq 0$. Moreover, $\varphi(r)=\varphi(R_{1})$ and hence $\Phi(r)=\Phi(R_{1})+(r-R_{1})\varphi(R_{1})$ for $r\geq R_{1}$. Now, we define the constant $c>0$ to be
$c:=\dfrac{D^{2}}{4\int_{0}^{R_{2}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s}$ (3.2)
Next, we consider the function
$g(r)=1-\frac{2c}{D^{2}}\int_{0}^{r}\Phi(s)\varphi(s)^{-1}\mathrm{d}s,\ \ \forall r\geq 0.$
Note that $\frac{1}{2}\leq g(r\wedge R_{2})\leq 1$. Finally, the function $f:[0,\infty)\to[0,\infty)$ is given by
$f(r)=\int_{0}^{r}\varphi(s)g(s\wedge R_{2})\mathrm{d}s,\ \ \forall r\geq 0.$ (3.3)
Note that $f(0)=0$, $f(r)>0$ for $r>0$, and $f\in C^{1}(0,\infty)\cap C^{2}((0,\infty)\setminus\{R_{2}\})$ with
$f^{\prime}(r)=\varphi(r)g(r\wedge R_{2})\in(0,1],$ (3.4)
$\begin{aligned} f^{\prime\prime}(r)&=-\frac{2f^{\prime}(r)}{D^{2}}(r\kappa^{*}(r)+A)-\frac{2c}{D^{2}}\Phi(r)I_{\{r<R_{2}\}}\\ &\leq-\frac{2f^{\prime}(r)}{D^{2}}(r\kappa^{*}(r)+A)-\frac{2c}{D^{2}}f(r)I_{\{r<R_{2}\}}\leq 0.\end{aligned}$ (3.5)
Consequently, $f$ is a strictly increasing and concave function. Note also that $f$ is linear on $R_{2},\infty)$.
Finally, it is clear that f(|xy|)=f(|yx|)f(|x-y|)=f(|y-x|) and $f(|x-y|)\leq f(|x-z|)+f(|z-y|)$ for all $x,y,z\in\mathbb{R}^{d}$. Thus,
$\rho(x,y):=f(|x-y|)$ (3.6)
is a well defined metric on $\mathbb{R}^{d}$.
The following lemma gives linear bounds on the function $f$.
Lemma 3.2.
For the functions $f$, $\varphi$, and $\Phi$ defined above, we have
$r\varphi(R_{1})\leq\Phi(r)\leq\frac{D^{2}f(r)}{2}\leq\frac{D^{2}\Phi(r)}{2}\leq\frac{D^{2}r}{2}.$
Proof.
The first inequality follows since $\varphi(r)$ is decreasing for $r<R_{1}$ and $\varphi(r)=\varphi(R_{1})$ for $r\geq R_{1}$, so
$r\varphi(R_{1})-\Phi(r)=\int_{0}^{r}(\varphi(R_{1})-\phi(s))\mathrm{d}s\leq 0.$
Next, the observation $2/D^{2}\leq g(r\wedge R_{2})\leq 1$ gives the second and the third inequalities. The last inequality follows since $\varphi(r)\leq 1$. ∎
Theorem 3.3.
Suppose Assumptions 1.1, 1.2, 1.3, 1.4, 1.5, and 3.1 hold. Let $\mu_{t}$ and $\nu_{t}$ denote the marginal laws of a strong solution $(X_{t})$ of equation (1.1) with initial conditions $X_{0}\sim\mu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d})$ and $X_{0}\sim\nu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d})$, respectively. Then there exist constants $C$ and $\gamma$ such that
$\mathcal{W}_{\rho}(\mu_{t},\nu_{t})$ $\leq e^{-\gamma t}\mathcal{W}_{\rho}(\mu_{0},\nu_{0})$ (3.7)
$\mathcal{W}_{1}(\mu_{t},\nu_{t})$ $\leq Ce^{-\gamma t}\mathcal{W}_{1}(\mu_{0},\nu_{0}),$ (3.8)
where the metric $\rho$ is given in (3.6) and the constants are given by
$\gamma=c-\frac{L_{1}D^{2}}{2\varphi(R_{1})}\ \ \ \text{ and }\ \ \ C=\frac{D^{2}}{2\varphi(R_{1})},$ (3.9)
in which $c$, $R_{1}$, $D$ and $\varphi$ are defined as above.
Remark 3.4.
In application, one would hope for $\gamma>0$. This requires $L_{1},$ to be small when $R_{1}$ and $R_{2}$ are large. Such requirements are, in fact, natural. We notice that $\kappa(r)<0$ is needed to ensure recurrence of the radial process inside some ball of radius $R_{2}$. However, if $L_{1}$ is large, one could have multiple distinct stationary solutions to (1.1) inside that ball.
Proof.
We consider the function $f$ defined in (3.3) and note that using the coupling $U_{t}=(X_{t},Y_{t})$ from Section 2, in which we assume without loss of generality that $\mathcal{W}_{\rho}(\mu_{0},\nu_{0})=\mathbb{E}[\rho(X_{0},Y_{0})]=\mathbb{E}[f(r_{0})]$, we have
$\mathcal{W}_{\rho}(\mu_{t},\nu_{t})\leq\mathbb{E}[\rho(X_{t},Y_{t})]=\mathbb{E}[f(|X_{t}-Y_{t}|)]=\mathbb{E}[f(r_{t})].$
Therefore, we need to estimate $\mathbb{E}[f(r_{t})]$, in which $r_{t}$ satisfies (2.1). Applying the Itô-Tanaka formula (see, for example, (Revuz and Yor, , 1999, Theorem 22.5)) to $f(r_{t})$, we get almost surely
$f(r_{t})=$ $\ f(r_{0})+\int_{0}^{t}f^{\prime}(r_{s})\mathrm{d}r_{s}+\frac{1}{2}\int_{-\infty}^{\infty}L_{s}^{x}\mu_{f}(\mathrm{d}x)$
$=$ $\ f(r_{0})+\int_{0}^{t}I_{\{r_{s}>0\}}f^{\prime}(r_{s})\langle e_{s},\beta_{s}\rangle\mathrm{d}s+\int_{0}^{t}I_{\{r_{s}>0\}}f^{\prime}(r_{s})\frac{\mathrm{rc}^{2}(U_{s})}{2r_{s}}\left(\mathrm{tr}(\alpha_{s}\alpha_{s}^{\top})-|\alpha_{s}^{\top}e_{s}|^{2}\right)\mathrm{d}s$
$+\int_{0}^{t}I_{\{r_{s}>0\}}f^{\prime}(r_{s})\frac{\mathrm{sc}^{2}(U_{s})}{2r_{s}}\left(\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2}\right)\mathrm{d}s$ (3.10)
$+\int_{0}^{t}I_{\{r_{s}>0\}}f^{\prime}(r_{s})[\mathrm{rc}(U_{s})\langle e_{s},\alpha_{s}\mathrm{d}B_{s}^{2}\rangle+\mathrm{sc}(U_{s})\langle e_{s},\Delta_{s}\mathrm{d}B_{s}^{1}\rangle]+\frac{1}{2}\int_{-\infty}^{\infty}L_{t}^{x}\mu_{f}(\mathrm{d}x),$
where $L_{t}^{x}$ is the right-continuous local time of $(r_{t})$, and $\mu_{f}$ is the nonpositive measure representing the second derivative of $f$. Furthermore, from the Itô-Tanaka formula we have for every measurable function $g:\mathbb{R}\to\mathbb{R}_{+}$
$\int_{0}^{t}g(r_{s})\mathrm{d}[r]_{s}=\int_{-\infty}^{\infty}g(x)L_{t}^{x}\mathrm{d}x.$
Thanks to (1.3) and (1.4), we can use the same argument as that in (Lindvall and Rogers, , 1986, section 3) to show that
$|\alpha_{s}^{\top}e_{s}|^{2}\geq D^{2}>0.$ (3.11)
Since $\delta<1\leq R_{1}$, it follows that $r_{s}\geq\delta$ if $I_{\{r_{s}\in\{R_{1},R_{2}\}\}}=1$. Moreover, on the set $\{r_{s}\geq\delta\}$, $Z_{s}\neq 0$, and we have $\mathrm{sc}^{2}(U_{s})=0$ and $\mathrm{rc}^{2}(U_{s})=1$ therefore,
$\mathrm{d}[r]_{s}=I_{\{Z_{s}\neq 0\}}[\mathrm{rc}^{2}(U_{s})|\alpha_{s}^{\top}e_{s}|^{2}+\mathrm{sc}^{2}(U_{s})|\Delta_{s}^{\top}e_{s}|^{2}]\mathrm{d}s=|\alpha_{s}^{\top}e_{s}|^{2}\mathrm{d}s\geq D^{2}\mathrm{d}s.$
Consequently, we have
$\text{Leb}\{0\leq s\leq t:r_{s}\in\{R_{1},R_{2}\}\}$ $=\int_{0}^{t}I_{\{r_{s}\in\{R_{1},R_{2}\}\}}\mathrm{d}s\leq\frac{1}{D^{2}}\int_{0}^{t}I_{\{r_{s}\in\{R_{1},R_{2}\}\}}\mathrm{d}[r]_{s}$
$=\frac{1}{D^{2}}\int_{\{R_{1},R_{2}\}}L_{t}^{x}\mathrm{d}x=0\ \ \ \text{a.s.}$
where $\text{Leb}\{0\leq s\leq t:r_{s}\in\{R_{1},R_{2}\}\}$ is the Lebesgue measure. Furthermore, recall that $f$ is concave and that $f\in C^{1}(0,\infty)\cap C^{2}((0,\infty)\setminus\{R_{1},R_{2}\})$. Hence,
$\int_{-\infty}^{\infty}L_{t}^{x}\mu_{f}(\mathrm{d}x)$ $\leq\int_{0}^{t}f^{\prime\prime}(r_{s})\mathrm{d}[r]_{s}$ (3.12)
$=\int_{0}^{t}f^{\prime\prime}(r_{s})\left(\mathrm{rc}^{2}(U_{s})|\alpha_{s}^{\top}e_{s}|^{2}+\mathrm{sc}^{2}(U_{s})|\Delta_{s}^{\top}e_{s}|^{2}\right)\mathrm{d}s\ \text{ a.s.}$
We now plug (3.12) into (3.1) to obtain
$\mathrm{d}f(r_{t})$
$\leq I_{\{r_{t}>0\}}\bigg[f^{\prime}(r_{t})\bigg(\!\langle e_{t},\beta_{t}\rangle+\frac{\mathrm{rc}^{2}(U_{t})}{2r_{t}}(\mathrm{tr}(\alpha_{t}\alpha_{t}^{\top})-|\alpha_{t}^{\top}e_{t}|^{2})+\frac{\mathrm{sc}^{2}(U_{t})}{2r_{t}}(\mathrm{tr}(\Delta_{t}\Delta_{t}^{\top})-|\Delta_{t}^{\top}e_{t}|^{2})\bigg)\bigg]\mathrm{d}t$
$\quad+I_{\{r_{t}>0\}}\frac{f^{\prime\prime}(r_{t})}{2}\left(\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}+\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|^{2}\right)\mathrm{d}t$ (3.13)
$\quad+I_{\{r_{t}>0\}}f^{\prime}(r_{t})\left(\mathrm{rc}^{2}(U_{t})\langle e_{t},\alpha_{t}\mathrm{d}B^{2}_{t}\rangle+\mathrm{sc}^{2}(U_{t})\langle e_{t},\Delta_{t}\mathrm{d}B^{1}_{t}\rangle\right).$
Using integration by parts and (3.1), we have for the constant $\gamma=c-\frac{L_{1}D^{2}}{2\varphi(R_{1})}$,
$\mathbb{E}[e^{\gamma t}f(r_{t})-f(r_{0})]\leq\int_{0}^{t}e^{\gamma s}\mathbb{E}[\mathcal{E}_{s}]\mathrm{d}s,$ (3.14)
where
$\mathcal{E}_{s}:=$ $\,\gamma f(r_{s})+f^{\prime}(r_{s})I_{\{r_{s}>0\}}\langle e_{s},\beta_{s}\rangle+\frac{f^{\prime\prime}(r_{s})}{2}I_{\{r_{s}>0\}}(\mathrm{rc}^{2}(U_{s})|\alpha_{s}^{\top}e_{s}|^{2}+\mathrm{sc}^{2}(U_{s})|\Delta_{s}^{\top}e_{s}|^{2})$
$+I_{\{r_{s}>0\}}\frac{f^{\prime}(r_{s})}{2r_{s}}\big[\mathrm{rc}^{2}(U_{s})\left(\mathrm{tr}(\alpha_{s}\alpha_{s}^{\top})-|\alpha_{s}^{\top}e_{s}|^{2}\right)+\mathrm{sc}^{2}(U_{s})\left(\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2}\right)\big].$ (3.15)
We now attempt to bound $\mathcal{E}_{s}$ so that $\mathbb{E}[\mathcal{E}_{s}]\leq(\delta)$. To that end we will first find bounds for each term and substitute in the choice of $\gamma$. First, assumption 1.1 along with Lemma 3.2 and the fact that $\mathcal{W}_{1}(\mu_{s},\nu_{s})\leq\mathbb{E}[r_{s}]$ yields the bound
$I_{\{r_{s}>0\}}\langle e_{s},\beta_{s}\rangle$ $\leq I_{\{r_{s}>0\}}(r_{s}\kappa(r_{s})+L_{1}\mathcal{W}_{1}(\mu_{s},\nu_{s}))$ (3.16)
$\leq I_{\{r_{s}\geq\delta\}}r_{s}\kappa(r_{s})+I_{\{r_{s}<\delta\}}\kappa_{\delta}\delta+\frac{L_{1}D^{2}}{2\varphi(R_{1})}\mathbb{E}[f(r_{s})],$
where $\kappa_{r}$ is define in (1.5). Since $\mathrm{sc}^{2}(U_{s})+\mathrm{rc}^{2}(U_{s})=1$ and $f^{\prime}(r_{s})\leq 1$, using Lemma 2.2, we get on the set $\{0<r_{s}<\delta\}$
$\frac{f^{\prime}(r_{s})}{2r_{s}}\big[\mathrm{rc}^{2}(U_{s})\left(\mathrm{tr}(\alpha_{s}\alpha_{s}^{\top})-|\alpha_{s}^{\top}e_{s}|^{2}\right)+\mathrm{sc}^{2}(U_{s})\left(\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2}\right)\big]$ (3.17)
$\ \ \leq\frac{f^{\prime}(r_{s})}{2r_{s}}\big[\mathrm{rc}^{2}(U_{s})\cdot 2dL_{2}r_{s}^{2}+\mathrm{sc}^{2}(U_{s})\cdot 2dL_{2}r_{s}^{2}\big]=dL_{2}f^{\prime}(r_{s})r_{s}\leq dL_{2}\delta.$
Also recall that $\mathrm{sc}(U_{s})=0$ and $\mathrm{rc}(U_{s})=1$ for $r_{s}\geq\delta$. Therefore, using (2.8) in Lemma 2.2 and the definition of $A$ in (2.10), we get on the set $\{r_{s}\geq\delta\}$,
$\frac{\mathrm{tr}(\alpha_{s}\alpha_{s}^{\top})-|\alpha_{s}^{\top}e_{s}|^{2}}{2r_{s}}=\frac{\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2}}{2r_{s}}\leq\frac{d\|\Delta_{s}\||r_{s}|^{2}-|\Delta^{\top}r_{s}|^{2}}{2r_{s}^{3}}\leq\frac{A}{2}.$ (3.18)
In the above we used the bound $\mathrm{tr}(\Delta\Delta)^{\top}\leq rank(\Delta)|\Delta|^{2}\leq d|\Delta|^{2}$. Now plugging (3.16), (3.17), and (3.18) into (3.1), we obtain
$\mathcal{E}_{s}\leq$ $\,\frac{L_{1}D^{2}}{2\varphi(R_{1})}(\mathbb{E}[f(r_{s})]-f(r_{s}))+I_{\{r_{s}<\delta\}}\left(cf(r_{s})+\kappa_{\delta}\delta+dL_{2}\delta\right)$
$+I_{\{r_{s}\geq\delta\}}\left(cf(r_{s})+f^{\prime}(r_{s})\left(r_{s}\kappa(r_{s})+\frac{\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2}}{2r_{s}}\right)+\frac{f^{\prime\prime}(r_{s})}{2}|\alpha_{s}^{\top}e_{s}|^{2}\right)$
$\leq$ $\,\frac{L_{1}D^{2}}{2\varphi(R_{1})}(\mathbb{E}[f(r_{s})]-f(r_{s}))+I_{\{r_{s}<\delta\}}(c+\kappa_{\delta}+dL_{2})\delta$ (3.19)
$+I_{\{r_{s}\geq\delta\}}\bigg(cf(r_{s})+f^{\prime}(r_{s})\bigg(r_{s}\kappa(r_{s})+\frac{A}{2}\bigg)+\frac{D^{2}}{2}f^{\prime\prime}(r_{s})\bigg),$
where we used (3.11), (3.18), to derive the last inequality.
Note that the first term on the right-hand side of (3.1) is zero in expectation. The second term goes to zero as $\delta\to 0$. Therefore if we can show the last term is nonpositive, we will have $\mathbb{E}[\mathcal{E}_{s}]\leq O(\delta)$. To this end, we have from (3.5) that
$I_{\{r_{s}<R_{2}\}}\bigg[cf(r_{s})+f^{\prime}(r_{s})\bigg(r_{s}\kappa(r_{s})+\frac{A}{2}\bigg)+\frac{D^{2}}{2}f^{\prime\prime}(r_{s})\bigg]$ (3.20)
$\ \ \leq I_{\{r_{s}<R_{2}\}}\bigg[cf(r_{s})+f^{\prime}(r_{s})(r_{s}\kappa^{*}(r_{s})+A)+\frac{D^{2}}{2}f^{\prime\prime}(r_{s})\bigg]\leq 0.$
Next, using the facts that $\varphi(s)=\varphi(R_{1})$ and $\Phi(s)=\Phi(R_{1})+\varphi(R_{1})(s-R_{1})$ for $s\geq R_{1}$, we can show
$\int_{0}^{R_{2}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s\geq\int_{R_{1}}^{R_{2}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s\geq\frac{\Phi(R_{2})(R_{2}-R_{1})}{2\varphi(R_{1})}.$ (3.21)
On the other hand, since $\Phi$ is increasing and concave with $\Phi(0)=0$, on the set $\{r_{s}\geq R_{2}\}$, we have $\frac{\Phi(r_{s})}{r_{s}}\leq\frac{\Phi(R_{2})}{R_{2}}$ and hence
$\frac{r_{s}}{R_{2}}\geq\frac{\Phi(r_{s})}{\Phi(R_{2})}.$ (3.22)
Finally, we note that for $r\geq R_{2}$, $f^{\prime\prime}(r)=0$ and $f^{\prime}(r)=\varphi(R_{1})g(R_{2})=\frac{2\varphi(R_{1})}{D^{2}}$. Now using the definition of $R_{2}$, we get on the set $\{r_{s}\geq R_{2}\}$
$f^{\prime}(r_{s})(r_{s}\kappa(r_{s})+A)$ $=\frac{2\varphi(R_{1})}{D^{2}}(r_{s}\kappa(r_{s})+A)\leq\frac{2\varphi(R_{1})}{D^{2}}\left(-r_{s}\left(\frac{D^{2}}{R_{2}(R_{2}-R_{1})}+\frac{A}{R_{2}}\right)+A\right)$
$\leq-\frac{2\varphi(R_{1})}{D^{2}}\frac{D^{2}r_{s}}{R_{2}(R_{2}-R_{1})}=\frac{-2r_{s}\varphi(R_{1})}{R_{2}(R_{2}-R_{1})}\leq\frac{-2\varphi(R_{1})\Phi(r_{s})}{\Phi(R_{2})(R_{2}-R_{1})}$
$\leq-\frac{\Phi(r_{s})}{\int_{0}^{R_{2}}\Phi(s)\varphi(s)^{-1}ds}\leq-\frac{4c}{D^{2}}\Phi(r_{s})\leq-\frac{4c}{D^{2}}f(r_{s}),$
where we used (3.22), followed by (3.21), followed by the definition of $c$ (3.2), and finally lemma 3.2 to derive the last four inequalities. This together with (3.5) imply that
$I_{\{r_{s}\geq R_{2}\}}\bigg[cf(r_{s})+f^{\prime}(r_{s})\bigg(r_{s}\kappa(r_{s})+\frac{A}{2}\bigg)+\frac{D^{2}}{2}f^{\prime\prime}(r_{s})\bigg]$ (3.23)
$\ \ \leq I_{\{r_{s}\geq R_{2}\}}[cf(r_{s})+f^{\prime}(r_{s})(r_{s}\kappa(r_{s})+A)]\leq 0.$
Using (3.20) and (3.23) in (3.1) and then taking expectations, we obtain $\mathbb{E}[\mathcal{E}_{s}]\leq(c+\kappa_{\delta}+dL_{2})\delta$; which, together with (3.14), gives us
$\mathbb{E}[e^{\gamma t}f(r_{t})-f(r_{0})]\leq(c+\kappa_{\delta}+dL_{2})\delta\frac{e^{\gamma t}}{\gamma}.$
Then it follows that
$\mathcal{W}_{\rho}(\mu_{t},\nu_{t})$ $\leq\mathbb{E}[f(r_{t})]\leq e^{-\gamma t}\mathbb{E}[f(r_{0})]+\frac{(c+\kappa_{\delta}+dL_{2})\delta}{\gamma}$
$=e^{-\gamma t}\mathcal{W}_{\rho}(\mu_{0},\nu_{0})+\frac{(c+\kappa_{\delta}+dL_{2})\delta}{\gamma}.$
Recall that $\gamma$ is independent of the coupling, so it does not depend on $\delta$. Therefore, since $\delta>0$ can be chosen arbitrarily small, (3.7) follows.
Finally, Lemma 3.2 leads to
$\mathcal{W}_{1}(\mu_{t},\nu_{t})\leq\frac{D^{2}}{2\varphi(R_{1})}\mathbb{E}[f(r_{t})]\leq\frac{D^{2}}{2\varphi(R_{1})}e^{-\gamma t}\mathcal{W}_{\rho}(\mu_{0},\nu_{0})\leq\frac{D^{2}}{2\varphi(R_{1})}e^{-\gamma t}\mathcal{W}_{1}(\mu_{0},\nu_{0}),$ (3.24)
which shows (3.8), completing the proof. ∎
3.2 Contraction Result Under Dissipative Condition
We now replace Assumption 3.1 with the following assumption.
Assumption 3.5 (Drift).
There exist positive constants $R$ and $\lambda$ such that
$\langle x,b(x,\mu)\rangle\leq-\lambda|x|^{2}+L_{4}|x|\mu(|\cdot|),\quad\forall|x|\geq R.$
Theorem 3.6.
Assume assumptions 1.1, 1.2, 1.3, 1.4, 1.5, and 3.5 with $L_{4}\leq L_{1}\leq\lambda/2$. Let $\mu_{t}$ and $\nu_{t}$ denote the marginal laws of a strong solution $(X_{t})$ to (1.1) with initial laws $\mu_{0},\nu_{0}$, respectively. Then there exist a concave, bounded, nondecreasing continuous function $f:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $f(0)=0$, a Lyapunov function $V(x)$ and constants $c,\varepsilon,K_{1},K_{6}\in(0,\infty)$ such that
  1. (i)
    For any $K\in(0,\infty)$ there is an $L_{1}^{*}>0$ such that for any $L_{1}<L_{1}^{*}$ if the initial laws have $\mu_{0}(V),\nu_{0}(V)\leq K$, then
    $\mathcal{W}_{\rho_{1}}(\mu_{t},\nu_{t})$ $\leq e^{-ct}\mathcal{W}_{\rho_{1}}(\mu_{0},\nu_{0}),\ \ \ \text{and}$ (3.25)
    $\mathcal{W}_{1}(\mu_{t},\nu_{t})$ $\leq K_{1}e^{-ct}\mathcal{W}_{\rho_{1}}(\mu_{0},\nu_{0}).$ (3.26)
  2. (ii)
    There is an $L_{1}^{**}>0$ such that for any $L_{1}<L_{1}^{**}$ and initial laws $\mu_{0},\nu_{0}.$
    $\mathcal{W}_{\rho_{1}}(\mu_{t},\nu_{t})\leq e^{-ct}(\mathcal{W}_{\rho_{1}}(\mu_{0},\nu_{0})+K_{6}[\varepsilon\mu_{0}(V)+\varepsilon\nu_{0}(V)]^{2}),$ (3.27)
where $c=\frac{1}{2}\min\{\lambda/2,\eta D^{2}/8\}$ with $\eta$ given below in (3.2), and the semi-metric $\rho_{1}$ is given by
$\rho_{1}(x,y)=f(|x-y|)(1+\varepsilon V(x)+\varepsilon V(y)).$
Remark 3.7.
Note again the Lipschitz coefficient for the law dependence, $L_{1}$, needs to be small enough. The choice of $c$ utilizes the rate, $\lambda$, at which the process returns to the ball, as well as the rate of convergence inside the ball due to the coupling.
We will construct a function $f$ similar to that from Section 3.1, but we will first need the following lemma.
Lemma 3.8.
Let $V(x)=1+|x|^{2}$. Suppose Assumptions 1.1, 1.2, 1.3, and 3.5 hold, and $L_{4}\leq L_{1}\leq\frac{\lambda}{2}$. Then there exists a constant $L\in(0,\infty)$ such that
$\mathrm{d}V(X_{t})\leq(L-\lambda V(X_{t})-\lambda|X_{t}|^{2}+2L_{1}|X_{t}|\mathbb{E}[|X_{t}|])\mathrm{d}t+2\langle X_{t},\sigma(X_{t})\mathrm{d}B_{t}\rangle,$ (3.28)
and
$\mathbb{E}[V(X_{t})]\leq\frac{L}{\lambda}+e^{-\lambda t}\mathbb{E}[V(X_{0})].$ (3.29)
Proof.
By Itô’s formula with Assumptions 1.4 and 3.5 and the fact that $L_{4}\leq L_{1}$, we have
$\mathrm{d}V(X_{t})$ $=\left(2\langle X_{t},b(X_{t},\mu_{t})\rangle+\mathrm{tr}(\sigma^{\top}(X_{t})\sigma(X_{t}))\right)\mathrm{d}t+2\langle X_{t},\sigma(X_{t})\mathrm{d}B_{t}\rangle$
$\leq I_{\{|X_{t}|\leq R\}}2\langle X_{t},b(X_{t},\mu_{t})\rangle\mathrm{d}t+I_{\{|X_{t}|\geq R\}}2(-\lambda|X_{t}|^{2}+L_{1}|X_{t}|\mathbb{E}[|X_{t}|])\mathrm{d}t$
$\quad+dM\mathrm{d}t+2\langle X_{t},\sigma(X_{t})\mathrm{d}B_{t}\rangle.$
We next recall the definition of $\kappa_{R}$ from (1.5) and use Assumption 1.1 to compute
$I_{\{|X_{t}|\leq R\}}\langle X_{t},b(X_{t},\mu_{t})\rangle$ $=I_{\{|X_{t}|\leq R\}}[\langle X_{t},b(0,\delta_{0})\rangle+\langle X_{t},b(X_{t},\mu_{t})-b(0,\delta_{0})\rangle]$
$\leq I_{\{|X_{t}|\leq R\}}[\langle X_{t},b(0,\delta_{0})\rangle+\kappa(|X_{t}|)|X_{t}|^{2}+L_{1}|X_{t}|\mathcal{W}_{1}(\mu_{t},\delta_{0})]$
$\leq\kappa_{R}R^{2}+R|b(0,\delta_{0})|+L_{1}I_{\{|X_{t}|\leq R\}}|X_{t}|\mathbb{E}[|X_{t}|];$
note that Assumption 1.3 implies $|b(0,\delta_{0})|<\infty$. Then we have
$\mathrm{d}V(X_{t})$ $\leq(-2\lambda I_{\{|X_{t}|\geq R\}}|X_{t}|^{2}+2L_{1}|X_{t}|\mathbb{E}[|X_{t}|]+C_{1})\mathrm{d}t+2\langle X_{t},\sigma(X_{t})\mathrm{d}B_{t}\rangle$
$\leq(-2\lambda|X_{t}|^{2}+2L_{1}|X_{t}|\mathbb{E}[|X_{t}|]+C_{2})\mathrm{d}t+2\langle X_{t},\sigma(X_{t})\mathrm{d}B_{t}\rangle$
where $C_{1}=dL_{2}+\kappa_{R}R^{2}+R|b(0,\delta_{0})|<\infty$ and $C_{2}:=C_{1}+2\lambda R^{2}$. This establishes (3.28) for $L=C_{2}+\lambda$.
Since $L_{1}\leq\frac{\lambda}{2}$, we can use (3.28) and integration by parts to obtain
$\mathrm{d}(e^{\lambda t}V(X_{t}))$ $=\lambda e^{\lambda t}V(X_{t})\mathrm{d}t+e^{\lambda t}\mathrm{d}V(X_{t})$
$\leq e^{\lambda t}\left(L-\lambda|X_{t}|^{2}+\lambda|X_{t}|\mathbb{E}[|X_{t}|]\right)\mathrm{d}t+2\langle X_{t},\sigma(X_{t})\mathrm{d}B_{t}\rangle,$
which, in turn, leads to
$\mathbb{E}[e^{\lambda t}V(X_{t})-V(X_{0})]$ $\leq\int_{0}^{t}e^{\lambda s}\left(L-\lambda\mathbb{E}[|X_{s}|^{2}]+\lambda\mathbb{E}[X_{s}]^{2}\right)\mathrm{d}s$
$\leq\int_{0}^{t}e^{\lambda s}L\mathrm{d}s\leq\frac{L}{\lambda}e^{\lambda t}.$
This gives (3.29) and completes the proof. ∎
Now we construct a new function $f:\mathbb{R}_{+}\to\mathbb{R}_{+}$ as follows. Using the function $V(x)=1+|x|^{2}$ and the constants $\lambda,L$ in Assumption 3.5 and Lemma 3.8, we can define
$S_{3}$ $=\{(x,y):V(x)+V(y)\leq\frac{2L}{\lambda}\},$
$S_{4}$ $=\{(x,y):V(x)+V(y)\leq\frac{8L}{\lambda}\},$
$R_{3}$ $=\sup_{(x,y)\in S_{3}}|x-y|,$
$R_{4}$ $=\sup_{(x,y)\in S_{4}}|x-y|.$
As before, put z=xyz=x-y and $r=|z|$. Then we define the function $f:\mathbb{R}_{+}\to\mathbb{R}_{+}$ by
$f(r)=\int_{0}^{r\wedge R_{4}}\varphi(s)g(s)\mathrm{d}s,$ (3.30)
where
$\varphi(r)$ $=e^{-h(r)},$
$\Phi(r)$ $=\int_{0}^{r}\varphi(s)\mathrm{d}s,$
$h(r)$ $=\frac{2}{D^{2}}\int_{0}^{r}(s\kappa^{*}(s)+A)\mathrm{d}s+\frac{8\tilde{M}}{D}r,$
$g(r)$ $=1-\frac{\eta}{4}\int_{0}^{r\wedge R_{4}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s-\frac{\xi}{4}\int_{0}^{r\wedge R_{3}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s,$
and
$\eta^{-1}$ $=\int_{0}^{R_{4}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s,\ \ \ \xi^{-1}=\int_{0}^{R_{3}}\Phi(s)\varphi(s)^{-1}\mathrm{d}s,\quad\tilde{M}=2M+\|\sigma(0)\|.$ (3.31)
Note that $\|\sigma(x)+\sigma(y)\|\leq\tilde{M}$ for all $x,y$. In the above and in the remainder of this subsection, $A$ is defined as in (2.10), $\kappa^{*}(r):=\kappa(r)\vee A$, and $D:=2\Lambda^{-1}-\sqrt{2}M$ as in (1.4). Note $D>0$ by Assumption 1.5. We further define the semi-metric
$\rho_{1}(x,y)=f(|x-y|)(1+\varepsilon V(x)+\varepsilon V(y)),\quad x,y\in\mathbb{R}^{d},$ (3.32)
where $\varepsilon=\frac{\xi D^{2}}{16L}$.
Note that $\varphi(r)\leq 1$ and is decreasing and $1/2\leq g(r)\leq 1$. Therefore, it follows that for $r<R_{4}$
$r\varphi(R_{4})\leq\Phi(r)\leq 2f(r)\leq 2\Phi(r)\leq 2r$ (3.33)
Furthermore, straightforward computations reveal that for $r<R_{4}$, we have
$f^{\prime}(r)$ $=\varphi(r)g(r),$ (3.34)
and
$f^{\prime\prime}(r)$ $=-h^{\prime}(r)f^{\prime}(r)-\left(\frac{\eta}{4}\Phi(r)\varphi(r)^{-1}+\frac{\xi}{4}\Phi(r)\varphi(r)^{-1}I_{\{r<R_{3}\}}\right)\varphi(r)$
$=-f^{\prime}(r)\left(\frac{2}{D^{2}}(r\kappa^{*}(r)+A)+\frac{8\tilde{M}}{D}\right)-\frac{\eta}{4}\Phi(r)-\frac{\xi}{4}\Phi(r)I_{\{r<R_{3}\}}$
$\leq-f^{\prime}(r)\left(\frac{2}{D^{2}}(r\kappa(r)+A)+\frac{8\tilde{M}}{D}\right)-\frac{\eta}{4}f(r)-\frac{\xi}{4}f(r)I_{\{r<R_{3}\}},$ (3.35)
where the last inequality follows from (3.33) and the fact that $\kappa(r)\leq\kappa^{*}(r)$.
We are now ready to prove Theorem 3.6. The essence of the proof is to bound $\mathbb{E}[\rho_{1}(X_{t},Y_{t})]$, in which $(X_{t},Y_{t})$ is the coupling process constructed in (2.3) in Section 2 and $\rho_{1}$ is the semi-metric defined in (3.32). In view of the definition of $\rho_{1}$, we will estimate the terms $f(r_{t})$, $G(X_{t},Y_{t})$, and their cross variation separately, where the function $f$ is defined in (3.30) and $G(x,y):=1+\varepsilon V(x)+\varepsilon V(y)$. These estimations together with integration by parts will then give us the desired result.
Proof of Theorem 3.6.
As we explained earlier, the goal is to bound $\mathbb{E}[\rho_{1}(X_{t},Y_{t})]$, where $(X_{t},Y_{t})$ is the coupling process given in (2.3); note that an arbitrary $\delta\in(0,1)$ was used in the construction of the coupling process. The rest of the proof is divided into five steps.
Step 1. In this step, we calculate a bound on $df(r_{t})$. Using the coupling process $U_{t}=(X_{t},Y_{t})$ and the same notation as those given in Section 2, similar arguments as those for (3.1) reveal that for $r_{t}\in(0,R_{4})\setminus\{R_{3}\}$, we still have
$\mathrm{d}f(r_{t})$
$\leq I_{\{r_{t}>0\}}f^{\prime}(r_{t})\bigg(\langle e_{t},\beta_{t}\rangle+\frac{\mathrm{rc}^{2}(U_{t})}{2r_{t}}\left(\mathrm{tr}(\alpha_{t}\alpha_{t}^{\top})-|\alpha_{t}^{\top}e_{t}|^{2}\right)+\frac{\mathrm{sc}^{2}(U_{t})}{2r_{t}}(\mathrm{tr}(\Delta_{t}\Delta_{t}^{\top})-|\Delta_{t}^{\top}e_{t}|^{2})\bigg)\mathrm{d}t$
$\quad+I_{\{r_{t}>0\}}\frac{f^{\prime\prime}(r_{t})}{2}\left(\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}+\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|^{2}\right)\mathrm{d}t$
$\quad+I_{\{r_{t}>0\}}f^{\prime}(r_{t})\left(\mathrm{rc}(U_{t})\langle e_{t},\alpha_{t}\mathrm{d}B_{t}^{2}\rangle+\mathrm{sc}(U_{t})\langle e_{t},\Delta_{t}\mathrm{d}B_{t}^{1}\rangle\right).$
Since $f$ is constant for $r\geq R_{4}$, we have $df(r_{t})=0$ on the set $\{r_{t}\geq R_{4}\}$.
We now focus on the set $\{r_{t}<R_{4}\}$. First, the same observation for (3.16) leads to
$I_{\{r_{t}>0\}}\langle e_{t},\beta_{t}\rangle\leq I_{\{r_{t}\geq\delta\}}r_{t}\kappa(r_{t})+I_{\{r_{t}<\delta\}}\kappa_{\delta}\delta+L_{1}\mathbb{E}[r_{t}].$
This, together with (3.34), (3.35), and the fact that $0<f^{\prime}(r)\leq 1$, allows us to get
$I_{\{r_{t}<R_{4}\}}\mathrm{d}f(r_{t})\leq$ $\,I_{\{r_{t}>0\}}\bigg[f^{\prime}(r_{t})\left(I_{\{r_{t}<\delta\}}\kappa_{\delta}\delta+I_{\{r_{t}\geq\delta\}}r_{t}\kappa(r_{t})+L_{1}\mathbb{E}[r_{t}]\right)\mathrm{d}t$
$+\frac{f^{\prime}(r_{t})}{2r_{t}}[\mathrm{sc}^{2}(U_{t})\left(\mathrm{tr}(\Delta_{t}^{\top}\Delta_{t})-|\Delta_{t}^{\top}e_{t}|^{2}\right)+\mathrm{rc}^{2}(U_{t})\left(\mathrm{tr}(\alpha_{t}^{\top}\alpha_{t})-|\alpha_{t}^{\top}e_{t}|^{2}\right)]\mathrm{d}t$
$-\frac{f^{\prime}(r_{t})}{2}\left(\frac{2}{D^{2}}(r_{t}\kappa(r_{t})+A)+\frac{8\tilde{M}}{D}\right)\left(\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}+\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|^{2}\right)\mathrm{d}t$
$-\left(\frac{\eta}{8}+\frac{\xi}{8}I_{\{r_{t}<R_{3}\}}\right)f(r_{t})\left(\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}+\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|^{2}\right)\mathrm{d}t$
$+f^{\prime}(r_{t})\left(\mathrm{rc}(U_{t})\langle e_{t},\alpha_{t}\mathrm{d}B_{t}^{2}\rangle+\mathrm{sc}(U_{t})\langle e_{t},\Delta_{t}\mathrm{d}B_{t}^{1}\rangle\right)\bigg]$
$\leq$ $\,\left(L_{1}f^{\prime}(r_{t})\mathbb{E}[r_{t}]+I_{\{0<r_{t}<\delta\}}(\kappa_{\delta}+K_{0})\delta\right)\mathrm{d}t$
$+I_{\{r_{t}\geq\delta\}}f^{\prime}(r_{t})\left[r_{t}\kappa(r_{t})+\frac{\mathrm{tr}(\Delta_{t}^{\top}\Delta_{t})-|\Delta_{t}^{\top}e_{t}|^{2}}{2r_{t}}-\frac{|\alpha_{t}^{\top}e_{t}|^{2}}{D^{2}}\left(r_{t}\kappa(r_{t})+A\right)\right]\mathrm{d}t$
$-I_{\{r_{t}\geq\delta\}}\frac{1}{8}|\alpha_{t}^{\top}e_{t}|^{2}f(r_{t})\left(\eta+\xi I_{\{r_{t}<R_{3}\}}\right)\mathrm{d}t$
$-\frac{8\tilde{M}f^{\prime}(r_{t})}{2D}I_{\{r_{t}>0\}}\left(\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}+\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|^{2}\right)\mathrm{d}t$
$+I_{\{r_{t}>0\}}f^{\prime}(r_{t})\left(\mathrm{rc}(U_{t})\langle e_{t},\alpha_{t}\mathrm{d}B_{t}^{2}\rangle+\mathrm{sc}(U_{t})\langle e_{t},\Delta_{t}\mathrm{d}B_{t}^{1}\rangle\right),$ (3.36)
where again $\kappa_{\delta}:=\sup\{|\kappa(r)|:r\in[0,\delta]\}$. Note that Lemma 2.2 is used to derive the last inequality with $K_{0}=dL_{2}$ as in (3.17). Also recall the bounds $|\alpha_{t}^{\top}e_{t}|^{2}\geq D^{2}$ and $f(r)\leq r$ derived in (3.11) and (3.33), respectively. Then we can estimate the second and fourth terms of (3.2) by
$I_{\{0<r_{t}<\delta\}}(\kappa_{\delta}+K_{0})\delta-I_{\{r_{t}\geq\delta\}}\frac{1}{8}|\alpha_{t}^{\top}e_{t}|^{2}f(r_{t})\left(\eta+\xi I_{\{r_{t}<R_{3}\}}\right)$
$\leq I_{\{0<r_{t}<\delta\}}(\kappa_{\delta}+K_{0})\delta-I_{\{r_{t}\geq\delta\}}\frac{D^{2}}{8}f(r_{t})\left(\eta+\xi I_{\{r_{t}<R_{3}\}}\right)$
$=I_{\{0<r_{t}<\delta\}}\left((\kappa_{\delta}+K_{0})\delta+\frac{D^{2}}{8}f(r_{t})\left(\eta+\xi I_{\{r_{t}<R_{3}\}}\right)\right)-\frac{D^{2}}{8}f(r_{t})\left(\eta+\xi I_{\{r_{t}<R_{3}\}}\right)$
$\leq I_{\{0<r_{t}<\delta\}}\left((\kappa_{\delta}+K_{0})\delta+\frac{D^{2}}{8}r_{t}\left(\eta+\xi I_{\{r_{t}<R_{3}\}}\right)\right)-\frac{D^{2}}{8}f(r_{t})\left(\eta+\xi I_{\{r_{t}<R_{3}\}}\right)$
$\leq\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}\right)\delta-\frac{D^{2}}{8}f(r_{t})\left(\eta+\xi I_{\{r_{t}<R_{3}\}}\right).$ (3.37)
Now, using the definition of $A$ in (2.10) and the bound (3.11), we notice the third term in (3.2) is
$f^{\prime}(r_{t})\left(r_{t}\kappa(r_{t})+\frac{\mathrm{tr}(\Delta_{t}^{\top}\Delta_{t})-|\Delta_{t}^{\top}e_{t}|^{2}}{2r_{t}}\right)-\frac{f^{\prime}(r_{t})|\alpha_{t}^{\top}e_{t}|^{2}}{D^{2}}\left(r_{t}\kappa(r_{t})+A\right)\leq 0.$ (3.38)
Plugging (3.2) and (3.38) into (3.2), we have for $r_{t}<R_{4}$,
$\mathrm{d}f(r_{t})\leq$ $\,-\frac{D^{2}}{8}f(r_{t})\left(\eta+\xi I_{\{r_{t}<R_{3}\}}\right)\mathrm{d}t+\left(L_{1}f^{\prime}(r_{t}))\mathbb{E}[r_{t}]+\delta\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}\right)\right)\mathrm{d}t$
$-I_{\{r_{t}>0\}}\frac{4\tilde{M}f^{\prime}(r_{t})}{D}\left(|\alpha_{t}^{\top}e_{t}|^{2}\mathrm{rc}^{2}(U_{t})+|\Delta_{t}^{\top}e_{t}|^{2}\mathrm{sc}^{2}(U_{t})\right)\mathrm{d}t$
$+I_{\{r_{t}>0\}}f^{\prime}(r_{t})\left(\mathrm{rc}^{2}(U_{t})\langle e_{t},\alpha_{t}\mathrm{d}B_{t}^{2}\rangle+\mathrm{sc}^{2}(U_{t})\langle e_{t},\Delta_{t}\mathrm{d}B_{t}^{1}\rangle\right).$ (3.39)
Step 2: This step aims to derive an upper bound for $\mathrm{d}G(X_{t},Y_{t})$ in (3.2). Recall that $G(x,y)=1+\varepsilon V(x)+\varepsilon V(y)$. Now we use (3.28) in Lemma 3.8 combined with the facts that $2L_{1}\leq\lambda$ and $V(x)\geq 2|x|$ to get
$\mathrm{d}V(X_{t})$ $\leq(L-\lambda V(X_{t}))\mathrm{d}t+\lambda V(X_{t})\mathbb{E}[V(X_{t})]\mathrm{d}t+2\langle X_{t},\sigma(X_{t})\mathrm{d}B_{t}\rangle,$
$\mathrm{d}V(Y_{t})$ $\leq(L-\lambda V(Y_{t}))\mathrm{d}t+\lambda V(Y_{t})\mathbb{E}[V(Y_{t})]\mathrm{d}t+2\langle Y_{t},\sigma(Y_{t})\mathrm{d}\hat{B}_{t}\rangle$
where
$B_{t}$ $=\int_{0}^{t}\mathrm{rc}(U_{s})\mathrm{d}B_{s}^{1}+\int_{0}^{t}\mathrm{sc}(U_{s})\mathrm{d}B_{s}^{2},\ \ \ \ \ \text{and}$
$\hat{B}_{t}$ $=\int_{0}^{t}\mathrm{rc}(U_{s})H_{s}\mathrm{d}B_{s}^{1}+\int_{0}^{t}\mathrm{sc}(U_{s})\mathrm{d}B_{s}^{2}.$
By the definition of $R_{1}$, on the set $\{r_{t}\geq R_{3}\}$, we have $V(X_{t})+V(Y_{t})\geq 2L/\lambda$ and hence $2L\varepsilon-\lambda\varepsilon V(X_{t})-\lambda\varepsilon V(Y_{t})\leq 0$. Similarly, on the set $\{r_{t}\geq R_{4}\}$, we have $V(X_{t})+V(Y_{t})\geq 8L/\lambda$ and so
$2L\varepsilon-\lambda\varepsilon V(X_{t})-\lambda\varepsilon V(Y_{t})$ $\leq 2L\varepsilon-\lambda\varepsilon\frac{4L}{\lambda}-\frac{\lambda\varepsilon}{2}V(X_{t})-\frac{\lambda\varepsilon}{2}V(Y_{t})$
$=-2L\varepsilon-\frac{\lambda\varepsilon}{2}V(X_{t})-\frac{\lambda\varepsilon}{2}V(Y_{t})$
$=-\frac{\lambda}{2}G(X_{t},Y_{t})+\frac{\lambda}{2}-2L\varepsilon$
$\leq-\min\left\{\frac{\lambda}{2},2L\varepsilon\right\}G(X_{t},Y_{t}).$
The last inequality follows because if $\lambda/2\leq 2L\varepsilon$, then $\frac{\lambda}{2}G(X_{t},Y_{t})+\frac{\lambda}{2}-2L\varepsilon\leq-\frac{\lambda}{2}G(X_{t},Y_{t})$, and if $2L\varepsilon\leq\lambda/2$, then
$-\frac{\lambda}{2}G(X_{t},Y_{t})+\frac{\lambda}{2}-2L\varepsilon=-2L\varepsilon G(X_{t},Y_{t})+\bigg(\frac{\lambda}{2}-2L\varepsilon\bigg)(1-G(X_{t},Y_{t}))\leq-2L\varepsilon G(X_{t},Y_{t}).$
Furthermore, since $\varepsilon=\frac{\xi D^{2}}{16L}$ and $\eta\leq\xi$, we have $2L\varepsilon\geq\frac{\eta D^{2}}{8}$ and therefore,
$2\varepsilon L-\varepsilon\lambda V(X_{t})-\varepsilon\lambda V(Y_{t})\leq-\min\left\{\frac{\lambda}{2},\frac{\eta D^{2}}{8}\right\}G(X_{t},Y_{t}).$
Then by Lemma 3.8,
$\mathrm{d}G(X_{t},Y_{t})=$ $\,\varepsilon\mathrm{d}V(X_{t})+\varepsilon\mathrm{d}V(Y_{t})$
$\leq$ $\left(2\varepsilon L-\varepsilon\lambda V(X_{t})-\varepsilon\lambda V(Y_{t})+2\varepsilon L_{1}(V(X_{t})\mathbb{E}[V(X_{t})]+V(Y_{t})\mathbb{E}[V(Y_{t})])\right)\mathrm{d}t$
$+2\varepsilon(\langle X_{t},\sigma(X_{t})\mathrm{d}B_{t}\rangle+\langle Y_{t},\sigma(Y_{t})\mathrm{d}\hat{B}_{t}\rangle)$
$\leq$ $\left(I_{\{r_{t}<R_{3}\}}2L\varepsilon-I_{\{r_{t}\geq R_{4}\}}\min\left\{\frac{\lambda}{2},\frac{\eta D^{2}}{8}\right\}G(X_{t},Y_{t})\right)\mathrm{d}t$ (3.40)
$+2\varepsilon L_{1}(V(X_{t})\mathbb{E}[V(X_{t})]+V(Y_{t})\mathbb{E}[V(Y_{t})])\mathrm{d}t$
$+2\varepsilon(\langle X_{t},\sigma(X_{t})\mathrm{d}B_{t}\rangle+\langle Y_{t},\sigma(Y_{t})\mathrm{d}\hat{B}_{t}\rangle.$
Step 3: We now estimate the cross variation of $f(r_{t})$ and $G(X_{t},Y_{t})$. The cross variation of $f(r_{t})$ and $G(X_{t},Y_{t})$ is given by
$\mathrm{d}[f(r),G(X,Y)]_{t}$ $=2\varepsilon f^{\prime}(r_{t})\mathrm{sc}^{2}(U_{t})\langle\Delta_{t}^{\top}e_{t},\sigma(X_{t})^{\top}X_{t}+\sigma(Y_{t})^{\top}Y_{t}\rangle\mathrm{d}t$ (3.41)
$\quad+2\varepsilon f^{\prime}(r_{t})\mathrm{rc}^{2}(U_{t})\langle\alpha_{t}^{\top}e_{t},\sigma(X_{t})^{\top}X_{t}+(\sigma(Y_{t})H_{t})^{\top}Y_{t}\rangle\mathrm{d}t.$
Furthermore, we notice that using the Cauchy-Schwarz inequality, followed by the bound (1.3) and the fact that $\|H\|=1$, we get
$2\varepsilon$ $f^{\prime}(r_{t})\mathrm{rc}^{2}(U_{t})\langle\alpha_{t}^{\top}e_{t},\sigma(X_{t})^{\top}X_{t}+(\sigma(Y_{t})H_{t})^{\top}Y_{t}\rangle$ (3.42)
$\leq 2\varepsilon f^{\prime}(r_{t})\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|(|\sigma(X_{t})^{\top}X_{t}|+|H_{t}^{\top}\sigma(Y_{t})^{\top}Y_{t}|)$
$\leq 2\varepsilon f^{\prime}(r_{t})\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|\tilde{M}(|X_{t}|+|Y_{t}|)$
$\leq\frac{\tilde{M}f^{\prime}(r_{t})}{D}\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}G(X_{t},Y_{t}),$
where we also used (3.11) to derive the last inequality. Similarly, since $f^{\prime}(r_{t})\leq 1$ and $\mathrm{sc}^{2}(U_{t})=0$ for $r_{t}\geq\delta$, we have from Assumption 1.4 that
$2\varepsilon$ $f^{\prime}(r_{t})\mathrm{sc}^{2}(U_{t})\langle\Delta_{t}^{\top}e_{t},\sigma(X_{t})^{\top}X_{t}+\sigma(Y_{t})^{\top}Y_{t}\rangle$ (3.43)
$\leq\tilde{M}f^{\prime}(r_{t})\mathrm{sc}^{2}(U_{t})|\Delta_{t}^{\top}e_{t}|G(X_{t},Y_{t})$
$\leq\tilde{M}\sqrt{L_{2}}f^{\prime}(r_{t})\mathrm{sc}^{2}(U_{t})r_{t}G(X_{t},Y_{t})$
$\leq\tilde{M}\sqrt{L_{2}}\delta G(X_{t},Y_{t}).$
Using (3.42) and (3.43) in (3.41), we obtain
$\mathrm{d}[f(r),G(X,Y)]_{t}$ $\leq\frac{\tilde{M}f^{\prime}(r_{t})}{D}\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}G(X_{t},Y_{t})\mathrm{d}t+\tilde{M}\sqrt{L_{2}}\delta G(X_{t},Y_{t})\mathrm{d}t.$ (3.44)
Step 4: In this step, we derive upper bounds for $\mathrm{d}(\rho_{1}(X_{t},Y_{t}))$ and $\mathbb{E}[e^{ct}\rho_{1}(X_{t},Y_{t})]$, in which $c$ is a suitable constant to be determined later. Now we can use integration by parts to combine (3.2), (3.2), and (3.44), and recall that $f^{\prime}(r_{t})=0$ for $r_{t}>R_{4}$ to get
$\mathrm{d}(\rho_{1}(X_{t},Y_{t}))$ $=G(X_{t},Y_{t})\mathrm{d}f(r_{t})+f(r_{t})\mathrm{d}G(X_{t},Y_{t})+\mathrm{d}[f(r),G(X,Y)]_{t}$
$\leq G(X_{t},Y_{t})\left[-\frac{\eta D^{2}}{8}f(r_{t})I_{r_{t}<R_{4}}-\frac{\xi D^{2}}{8}f(r_{t})I_{r_{t}<R_{3}}+L_{1}\mathbb{E}[r_{t}]\right]\mathrm{d}t$
$\quad+G(X_{t},Y_{t})\left[\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+\tilde{M}\sqrt{L_{2}}\right)\delta\right]\mathrm{d}t$
$\quad+G(X_{t},Y_{t})\left[-\frac{4\tilde{M}f^{\prime}(r_{t})}{D}\left(|\alpha_{t}^{\top}e_{t}|^{2}\mathrm{rc}^{2}(U_{t})+|\Delta_{t}^{\top}e_{t}|^{2}\mathrm{sc}^{2}(U_{t})\right)\right]\mathrm{d}t$
$\quad+f(r_{t})\left[\frac{\xi D^{2}}{8}I_{\{r_{t}<R_{3}\}}-\min\left\{\frac{\lambda}{2},\frac{\eta D^{2}}{8}\right\}G(X_{t},Y_{t})I_{\{r_{t}\geq R_{4}\}}\right]\mathrm{d}t$
$\quad+2f(r_{t})L_{1}\varepsilon(V(X_{t})\mathbb{E}[V(X_{t})]+V(Y_{t})\mathbb{E}[V(Y_{t})])\mathrm{d}t$
$\quad+\frac{\tilde{M}f^{\prime}(r_{t})}{D}\mathrm{rc}^{2}(U_{t})|\alpha_{t}^{\top}e_{t}|^{2}G(X_{t},Y_{t})\mathrm{d}t+\mathrm{d}M_{t}$
$\leq\left[-\min\left\{\frac{\lambda}{2},\frac{\eta D^{2}}{8}\right\}f(r_{t})G(X_{t},Y_{t})+L_{1}\mathbb{E}[r_{t}]G(X_{t},Y_{t})\right]\mathrm{d}t$ (3.45)
$\quad+G(X_{t},Y_{t})\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+\tilde{M}\sqrt{L_{2}}\right)\delta\mathrm{d}t$
$\quad+2f(r_{t})L_{1}\varepsilon(V(X_{t})\mathbb{E}[V(X_{t})]+V(Y_{t})\mathbb{E}[V(Y_{t})])\mathrm{d}t+\mathrm{d}M_{t},$
where $M_{t}$ is a local martingale.
Next, we observe that there exists a constant $K_{1}>0$ such that
$|x-y|\leq K_{1}f(|x-y|)G(x,y)=K_{1}\rho_{1}(x,y),\quad\text{ for all }x,y\in\mathbb{R}^{d}.$ (3.46)
Indeed, when $|x-y|<R_{2}$, (3.46) follows from (3.33) and the fact that $G(x,y)\geq 1$. When $|x-y|\geq R_{2}$, we have $f(|x-y|)=f(R_{2})$ and hence
$|x-y|\leq|x|+|y|\leq\frac{1}{2}(V(x)+V(y))\leq\frac{1}{2\varepsilon f(R_{2})}f(|x-y|)G(x,y).$
In view of (3.46), we have the bound
$L_{1}\mathbb{E}[r_{t}]G(X_{t},Y_{t})\leq L_{1}K_{1}\mathbb{E}[\rho_{1}(X_{t},Y_{t})]G(X_{t},Y_{t}).$ (3.47)
Furthermore, we have
$\varepsilon\mathbb{E}[V(X_{t})]V(X_{t})+\varepsilon\mathbb{E}[V(Y_{t})]V(Y_{t})\leq\varepsilon^{-1}G(X_{t},Y_{t})\mathbb{E}[G(X_{t},Y_{t})].$
Therefore
$2L_{1}\varepsilon f(r_{t})(\mathbb{E}[V(X_{t})]V(X_{t})+\mathbb{E}[V(Y_{t})]V(Y_{t}))\leq\frac{2L_{1}}{\varepsilon}f(r_{t})G(X_{t},Y_{t})\mathbb{E}[G(X_{t},Y_{t})].$ (3.48)
Now with $c=\frac{1}{2}\min\{\lambda/2,\eta D^{2}/8\}$ and using integration by parts and (3.2), (3.47), and (3.48), we get
$\mathbb{E}[e^{ct}\rho_{1}(X_{t},Y_{t})]\leq\rho_{1}(X_{0},Y_{0})+\int_{0}^{t}e^{cs}\mathbb{E}[\mathcal{J}_{s}]\mathrm{d}s,$ (3.49)
where
$\mathcal{J}_{s}=$ $-c\rho_{1}(X_{s},Y_{s})+L_{1}K_{1}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]G(X_{s},Y_{s})$ (3.50)
$+\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+\tilde{M}\sqrt{L_{2}}\right)\delta G(X_{s},Y_{s})+\frac{2L_{1}}{\varepsilon}\mathbb{E}[G(X_{s},Y_{s})]\rho_{1}(X_{s},Y_{s}).$
We first consider the third term in (3.50). From Lemma 3.8 and the fact that $c<\lambda$, we have
$\int_{0}^{t}e^{cs}\mathbb{E}[G(X_{s},Y_{s})]\mathrm{d}s$ $=\int_{0}^{t}e^{cs}\left(1+\varepsilon\mathbb{E}[V(X_{s})]+\varepsilon\mathbb{E}[V(Y_{s})]\right)\mathrm{d}s$
$\leq\int_{0}^{t}e^{cs}\left(1+\frac{2L}{\lambda}+\varepsilon e^{-\lambda s}(\mathbb{E}[V(X_{0})]+\mathbb{E}[V(Y_{0})])\right)\mathrm{d}s$
$\leq\frac{1}{c}\left(1+\frac{2L}{\lambda}\right)e^{ct}+\frac{\varepsilon}{\lambda-c}(\mathbb{E}[V(X_{0})+\mathbb{E}[V(Y_{0})]).$
Consequently, there exists a $K_{2}>0$ not depending on $\delta$ such that
$\left(\kappa_{\delta}+K_{0}+\frac{D^{2}(\eta+\xi)}{8}+\tilde{M}\sqrt{L_{2}}\right)\delta\int_{0}^{t}e^{cs}\mathbb{E}[G(X_{s},Y_{s})]\mathrm{d}s$
$\leq K_{2}\delta e^{ct}+K_{2}\varepsilon(\mathbb{E}[V(X_{0})+\mathbb{E}[V(Y_{0})])\delta.$
Now by Lemma 3.8, we have
$\mathbb{E}[G(X_{t},Y_{t})]\leq 1+\frac{2L\varepsilon}{\lambda}+e^{-\lambda t}\varepsilon(\mathbb{E}[V(X_{0})]+\mathbb{E}[V(Y_{0})]).$ (3.51)
Next we estimate the second and fourth terms in (3.50). Define $K_{3}=L_{1}K_{1}+2L_{1}/\varepsilon$, $K_{4}=1+2L\varepsilon/\lambda$, and $K_{5}=K_{5}(X_{0},Y_{0})=\varepsilon(\mathbb{E}[V(X_{0})]+\mathbb{E}[V(Y_{0})])$. Then we get
$L_{1}$ $K_{1}\int_{0}^{t}e^{cs}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathbb{E}[G(X_{s},Y_{s})]\mathrm{d}s+\frac{2L_{1}}{\varepsilon}\int_{0}^{t}e^{cs}\mathbb{E}[G(X_{s},Y_{s})]\mathbb{E}(\rho_{1}(X_{s},Y_{s})]\mathrm{d}s$
$=K_{3}\int_{0}^{t}e^{cs}\mathbb{E}[G(X_{s},Y_{s})]\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathrm{d}s$
$\leq K_{3}K_{4}\int_{0}^{t}e^{cs}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathrm{d}s+K_{3}K_{5}\int_{0}^{t}e^{(c-\lambda)s}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathrm{d}s.$
Then we have
$\int_{0}^{t}e^{cs}\mathbb{E}[\mathcal{J}_{s}]\mathrm{d}s\leq$ $\,(K_{2}e^{ct}+K_{2}K_{5})\delta+K_{3}K_{4}\int_{0}^{t}e^{cs}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathrm{d}s$
$+K_{3}K_{5}\int_{0}^{t}e^{(c-\lambda)s}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathrm{d}s-c\int_{0}^{t}e^{cs}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathrm{d}s.$
Plugging this into (3.49) gives us
$\mathbb{E}[e^{ct}\rho_{1}(X_{t},Y_{t})]$ $\leq\mathbb{E}[\rho_{1}(X_{0},Y_{0})]+(K_{2}e^{ct}+K_{2}K_{5})\delta$
$\quad+\left(K_{3}(K_{4}+K_{5})-c\right)\int_{0}^{t}e^{cs}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathrm{d}s.$
Step 5: In this final step, we establish (3.25), (3.26), and (3.27). First we recall that $c$ is independent of the coupling, and thus does not depend on $\delta$. We define
$L_{1}^{*}=(K_{1}+2/\varepsilon)^{-1}(K_{4}+K_{5})^{-1}c.$
If $L_{1}<L_{1}^{*}$, then $K_{3}(K_{4}+K_{5})\leq c$ and hence
$\mathbb{E}[\rho_{1}(X_{t},Y_{t})]\leq(K_{2}+K_{2}K_{5}e^{-ct})\delta+e^{-ct}\mathbb{E}[\rho_{1}(X_{0},Y_{0})].$
Letting $\delta\to 0$, we get (3.25). Furthermore, since $r_{t}\leq K_{1}\rho_{1}(X_{1},Y_{1})$, (3.26) follows immediately from (3.25).
Note that $K_{5}$ depends on the initial probability measures $\mu_{0},\nu_{0}$, so we only get a local result. To get (3.27) we define $L_{1}^{**}=(K_{1}+2/\varepsilon)^{-1}K_{4}^{-1}c$. Then, if $L_{1}<L_{1}^{**}$ so that $K_{3}K_{4}\leq c$, we have
$\mathbb{E}[\rho_{1}(X_{t},Y_{t})]\leq e^{-ct}\mathcal{W}_{\rho_{1}}(\mu_{0},\nu_{0})+(K_{2}+K_{2}K_{5}e^{-ct})\delta+e^{-ct}K_{3}K_{5}\int_{0}^{t}e^{(c-\lambda)s}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathrm{d}s.$
Note that (3.33) implies that $f(r)\leq f(R_{4})\leq R_{2}$ for all $r\geq 0$. Then, by equation (3.51), we have
$\int_{0}^{t}e^{(c-\lambda)s}\mathbb{E}[\rho_{1}(X_{s},Y_{s})]\mathrm{d}s\leq R_{4}\left(1+\frac{2L\varepsilon}{\lambda}+\varepsilon(\mathbb{E}[V(X_{0})]+\mathbb{E}[V(Y_{0})])\right)\int_{0}^{t}e^{(c-\lambda)s}\mathrm{d}s.$
Next, since $c=\frac{1}{2}\min\{\frac{\lambda}{2},\frac{D^{2}\eta}{8}\}<\lambda$ and $V(x)=1+|x|^{2}$, there exists a constant $K_{6}>0$ such that
$\mathbb{E}[\rho_{1}(X_{t},Y_{t})]$ $\leq e^{-ct}\mathcal{W}_{\rho_{1}}(\mu_{0},\nu_{0})+e^{-ct}K_{6}(\varepsilon\mathbb{E}[V(X_{0})]+\varepsilon\mathbb{E}[V(Y_{0})])^{2}$
$\ \ +(K_{2}+K_{2}\varepsilon(\mathbb{E}[V(X_{0})]+\mathbb{E}[V(Y_{0})])e^{-ct})\delta.$
Letting $\delta\to 0$ gives (3.27). ∎
4 Applications
In this section, we explore the applications of the contraction results presented in Section 3. We first show that under the conditions of Theorem 3.3, a propagation of chaos type result holds. Next, we show that under the conditions of Theorem 3.3 or 3.6, (1.1) possesses a unique invariant measure $\mu$ and that starting from any initial distribution $\nu_{0}$, the distribution of $X_{t}$ converges to $\mu$ at an exponential rate.
We need the following lemma, which shows that the solution to (1.1) with initial condition $X_{0}=0$ is $L^{1}$-bounded under certain conditions. The proof of the lemma is deferred to the Appendix in order not to disrupt the flow of presentation.
Lemma 4.1.
Suppose Assumptions 1.1, 1.2, 1.3, and 1.4 hold. Let $X_{t}(0)$ be the solution to (1.1) starting at $X_{0}=0$. If there exist $R_{0}>0$ and $K>0$ such that
$L_{1}+\kappa(r)<-K<0,\ \ \ \forall r>R_{0},$ (4.1)
then we have
$\sup_{t\geq 0}\mathbb{E}[|X_{t}(0)|]<\infty.$
4.1 Propagation of Chaos
We now show that under the same assumptions as those in Theorem 3.3, a propagation of chaos type result holds. To this end, let $(Y_{0}^{i},W^{i})_{i\geq 1}$ be independent copies of $(X_{0},W)$, in which $X_{0}$ is the initial condition of (1.1) with $\mathcal{L}_{X_{0}}=\mu_{0}$ and $W$ is a standard $d$-dimensional Brownian motion. For each $i=1,2,\dots,n$, we define the $n$-particle system by
$\mathrm{d}X_{t}^{n,i}=b(X_{t}^{n,i},\mu_{t}^{n})\mathrm{d}t+\sigma(X_{t}^{n,i})\mathrm{d}W_{t}^{i},\quad\ \ \ \mu_{t}^{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{t}^{i,n}},$ (4.2)
where $\delta_{X_{t}^{n,i}}$ is the Dirac measure and so $\mu_{t}^{n}$ is the empirical mean. We similarly define $n$-copies of the Mckean-Vlasov SDE (1.1) by
$\mathrm{d}Y_{t}^{i}=b(Y_{t}^{i},\mu_{t})\mathrm{d}t+\sigma(Y_{t})\mathrm{d}W_{t}^{i},\ \ \ \ \mu_{t}=\mathcal{L}(Y_{t}^{i}).$ (4.3)
For each $i$, we can define a coupling for $X_{t}^{n,i}$ of (4.2) and $Y_{t}^{i}$ of (4.3) similarly to that constructed in Section 2. In addition, we define $Z_{t}^{i}=X_{t}^{i,n}-Y_{t}^{i}$, and $r_{t}^{i}$, $e_{t}^{i}$, $U_{t}^{i}$, $\Delta_{t}^{i}$, $\beta_{t}^{i}$, and $\alpha_{t}^{i}$ in the same way as in Section 2.
Proposition 4.2.
Suppose the conditions of Theorem 3.3 and (4.1) hold. If the constant $\gamma$ given in (3.9) is positive, then the $n$-particle system (4.2) converges to (4.3) in the following sense:
$\mathcal{W}_{\rho}(\mu_{t}^{n},\mu_{t})\leq e^{-\gamma t}\mathcal{W}_{\rho}(\mu_{0}^{n},\mu_{0})+\frac{n(d)^{1/2}}{\gamma},$
where
$n(d)=C\begin{cases}n^{-1/2}\ \ \ \ \ &\text{ if }\ \ d<4,\\ n^{-1/2}\log n\ \ &\text{ if }\ \ d=4,\\ n^{-2/d}\ \ \ \ \ &\text{ if }\ \ d>4,\end{cases}$ (4.4)
with the positive constant $C$ depends on the dimension $d$ and the moment.
Proof.
First, we note that the process $Y_{t}^{i}$ is non-explosive for any initial value $Y_{0}^{i}$ with $\mathcal{L}(Y^{i}_{0})=\mu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Indeed using Lemma 3.2 and (3.24), we have
$\mathbb{E}[|Y_{t}^{i}-Y_{t}(0)|]\leq\frac{D^{2}}{2\varphi(R_{1})}\mathbb{E}[f(|Y_{t}(0)-Y_{t}^{i})|]\leq\frac{D^{2}}{2\varphi(R_{1})}e^{-\gamma t}\mathcal{W}_{\rho}(\delta_{0},\mu_{0}),$
where $Y_{t}(0)$ denotes the solution to (1.1) with initial condition $Y_{0}=0$. Thus, we have
$\sup_{t\geq 0}\mathbb{E}[|Y_{t}(0)-Y_{t}^{i}]\leq\frac{D^{2}}{2\varphi(R_{1})}\mathcal{W}_{\rho}(\delta_{0},\mu_{0})\leq\frac{D^{2}}{2\varphi(R_{1})}\mu_{0}(|\cdot|)<\infty.$
This, together with Lemma 4.1, implies that
$\sup_{t\geq 0}\mathbb{E}[|Y_{t}^{i}|]<\infty,\quad\forall i.$ (4.5)
From Lemma 2.1, we have
$\mathrm{d}r_{t}^{i}$ $=I_{\{r_{t}^{i}>0\}}\bigg[\frac{1}{2r_{t}^{i}}\big[\mathrm{rc}^{2}(U_{t}^{i})(\mathrm{tr}(\alpha_{t}^{i}\alpha_{t}^{i\top})-|\alpha_{t}^{i\top}e_{t}^{i}|^{2})+\mathrm{sc}^{2}(U_{t}^{i})(\mathrm{tr}(\Delta_{t}^{i}\Delta_{t}^{i\top})-|\Delta_{t}^{i\top}e_{t}^{i}|^{2})\big]\mathrm{d}t$
$\quad+\langle e_{t}^{i},\beta_{t}^{i}\rangle\mathrm{d}t+\mathrm{sc}(U_{t}^{i})\langle e_{t}^{i},\Delta_{t}\mathrm{d}B_{t}^{i,1}\rangle+\mathrm{rc}(U_{t}^{i})\langle e_{t}^{i},\alpha_{t}^{i}\mathrm{d}B_{t}^{i,2}\rangle\bigg].$
Furthermore, we can define $\mathcal{E}_{s}^{i}$ as in (3.1) to get
$\mathbb{E}[e^{\gamma t}f(r_{t}^{i})-f(r_{0}^{i})]\leq\int_{0}^{t}e^{\gamma s}\mathbb{E}[\mathcal{E}_{s}^{i}]\mathrm{d}s.$
Next, with $\nu_{t}^{n}:=\frac{1}{n}\sum_{k=1}^{n}\delta_{Y_{t}^{k}},$ we have the inequality
$\mathcal{W}_{1}(\mu_{t}^{n},\mu_{t})$ $\leq\mathcal{W}_{1}(\mu_{t},\nu_{t}^{n})+\mathcal{W}_{1}(\nu_{t}^{n},\mu_{t}^{n})\leq W_{1}(\mu_{t},\nu_{t}^{n})+\mathbb{E}[r_{t}^{i}].$
Substituting this into equation (3.16), we get
$I_{\{r_{s}^{i}>0\}}\langle e_{s}^{i},\beta^{i}_{s}\rangle$ $\leq I_{\{r_{s}^{i}>0\}}(r_{s}^{i}\kappa(r_{s}^{i})+L_{1}\mathcal{W}_{1}(\mu_{s}^{n},\mu_{s}))$
$\leq I_{\{r_{s}^{i}\geq\delta\}}r_{s}^{i}\kappa(r_{s}^{i})+I_{\{r_{s}^{i}<\delta\}}\kappa_{\delta}\delta+\frac{L_{1}D^{2}}{2\varphi(R_{1})}\mathbb{E}[f(r_{s}^{i})]+L_{1}\mathcal{W}_{1}(\mu_{s},\nu_{s}^{n}),$
Thus, with the above inequality, (3.1) becomes
$\mathcal{E}_{s}^{i}\leq$ $\ \frac{L_{1}D^{2}}{2\varphi(R_{1})}\left(\mathbb{E}[f(r_{s}^{i})]-f(r_{s}^{i})\right)+I_{\{r_{s}^{i}<\delta\}}(c+\kappa_{\delta}+dL_{2})\delta$
$+I_{\{r_{s}^{i}\geq\delta\}}\left(cf(r_{s}^{i})+f^{\prime}(r_{s}^{i})\left(r_{s}^{i}\kappa(r_{s}^{i})+\frac{A}{2}\right)+\frac{D}{2}f^{\prime\prime}(r_{s}^{i})\right)+L_{1}\mathcal{W}_{1}(\mu_{t},\nu_{t}^{i}).$
Therefore, following the remainder of the proof for Theorem 3.3, we get
$\mathbb{E}[e^{\gamma t}f(r_{t}^{i})]\leq\mathbb{E}[f(r_{0}^{i})]+\int_{0}^{t}e^{\gamma s}\mathbb{E}[\mathcal{W}_{1}(\mu_{t},\nu_{t}^{i})]\mathrm{d}s+(c+\kappa_{\delta}+dL_{2})\delta\frac{e^{\gamma t}}{\gamma}.$ (4.6)
Using Hölder’s inequality and (Carmona and Delarue, , 2018, Theorem 5.8), we obtain the bound
$\mathbb{E}[\mathcal{W}_{1}(\mu_{t},\nu_{t}^{i})]\leq\mathbb{E}[\mathcal{W}_{2}(\mu_{t},\nu_{t}^{i})]\leq\left(\mathbb{E}[\mathcal{W}_{2}^{2}(\mu_{t},\nu_{t}^{i})]\right)^{1/2}\leq n(d)^{1/2}.$
Thus we have
$\int_{0}^{t}e^{\gamma s}\mathbb{E}[\mathcal{W}_{1}(\mu_{t},\nu_{t}^{i})]\mathrm{d}s\leq\frac{e^{\gamma t}}{\gamma}n(d)^{1/2},$
which combined with (4.6) yields
$\mathcal{W}_{\rho}(\mu_{t}^{n},\mu_{t})\leq e^{-\gamma t}\mathcal{W}_{\rho}(\mu_{0}^{n},\mu_{0})+\frac{n(d)^{1/2}}{\gamma}+\frac{(c+\kappa_{\delta}+dL_{2})\delta}{\gamma}.$
Letting $\delta\to 0$ completes the proof. ∎
Proposition 4.3.
Suppose the conditions of Theorem 3.26 and (4.1) hold. If the constant $c$ given in (3.26) is positive, then the $n$-particle system (4.2) converges to (4.3) in the following sense:
$\mathcal{W}_{\rho_{1}}(\mu_{t}^{n},\mu_{t})\leq C(e^{-ct}\mathcal{W}_{\rho_{1}}(\mu_{0}^{n},\mu_{0})+\frac{n(d)^{1/2}}{c}),$
where $n(d)$ is defined as in (4.4) and some constant $C$.
Proof.
The result can be deduced using a proof that mirrors closely the one for Proposition 4.2, and defining instead $\mathcal{J}_{s}^{i}$ as in (3.50) and following the proof of Theorem 3.26. We note however that ∎
4.2 Exponential Ergodicity
As an application of the contraction results, we demonstrate that the McKean-Vlasov process given by (1.1) is exponentially ergodic under the conditions of Theorems 3.3 or 3.6. Let $(X_{t})_{t\geq 0}$ denote the solution to (1.1) with initial condition $X_{0}$. Suppose $\mathcal{L}(X_{0})=\mu$ and denote $P^{*}_{t}\mu=\mathcal{L}(X_{t})$ for $t\geq 0$. We have the semigroup property $P^{*}_{t+s}=P^{*}_{t}P^{*}_{s}$ for any $s,t\geq 0$. As usual, $\mu\in\mathcal{P}(\mathbb{R}^{d})$ is said to be an invariant measure for the semigroup $P_{\cdot}^{*}$ if $P^{*}_{t}\mu=\mu$ for any $t\geq 0$.
Theorem 4.4.
Suppose that (4.1) and the conditions of Theorem 3.3 hold with $\gamma>0$. Then $P_{t}^{*}$ admits an invariant measure $\mu\in\mathcal{P}_{1}(\mathbb{R}^{d})$ satisfying
$\mathcal{W}_{1}(P_{t}^{*}\nu_{0},\mu)\leq C\mathcal{W}_{1}(\mu,\nu_{0})e^{-\gamma t},\ \ \forall\nu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d}),$ (4.7)
where the constants $C$ and $\gamma$ are defined in Theorem 3.3.
Proof.
We adopt the shift coupling idea from the proof of (Wang, , 2018, Theorem 3.1). Let $\delta_{0}$ be the Dirac measure at 0. Then $P_{t}^{*}\delta_{0}=\mathcal{L}(X_{t})$. Next we define for any $s\geq 0$, $(\bar{X}_{t}:=X_{t+s}(0))_{t\geq 0}$. Note that $\bar{X}_{t}$ solves
$\mathrm{d}\bar{X}_{t}=b(\bar{X}_{t},\mathcal{L}(\bar{X}_{t})\mathrm{d}t+\sigma(\bar{X}_{t})\mathrm{d}\bar{W}_{t},\ \ \bar{X}_{0}=X_{s}(0)$
for the $d$-dimensional Brownian motion $\bar{W}_{t}=W_{t+s}-W_{s}$. Since (1.1) has a unique solution, we have
$P^{*}_{t}(P^{*}_{s}\delta_{0})=\mathcal{L}(\bar{X}_{t})=\mathcal{L}(X_{t+s}(0))=P^{*}_{t+s}\delta_{0},\ \ \forall s,t\geq 0.$ (4.8)
Now Theorem 3.3 and equation (4.8) combined with $\mathcal{L}(\bar{X}_{0})=P^{*}_{s}\delta_{0}$ gives us
$\mathcal{W}_{1}(P^{*}_{t+s}\delta_{0},P^{*}_{t}\delta_{0})$ $=\mathcal{W}_{1}(\mathcal{L}(X_{t}(P^{*}_{s}\delta_{0})),\mathcal{L}(X_{t}(0)))$
$\leq C\mathcal{W}_{1}(P^{*}_{s}\delta_{0},\delta_{0})e^{-\gamma t}$
$=Ce^{-\gamma t}\mathbb{E}[|X_{s}(0)|],\ \ \ s,t\geq 0$
Thanks to Lemma 4.1, $\sup_{s\geq 0}\mathbb{E}[|X_{s}(0)|]<\infty$. Therefore, we have
$\lim_{t\to\infty}\sup_{s\geq 0}\mathcal{W}_{1}(P^{*}_{t}\delta_{0},P^{*}_{t+s}\delta_{0})=0.$
Consequently, $\{P^{*}_{t}\delta_{0}\}_{t\geq 0}$ is Cauchy in $\mathcal{W}_{1}$. Thus, there exists some $\mu\in\mathcal{P}_{1}(\mathbb{R}^{d})$ such that
$\lim_{t\to\infty}\mathcal{W}_{1}(P^{*}_{t}\delta_{0},\mu)=0.$ (4.9)
Now, we combine (4.9) with Theorem 3.3 to derive
$\lim_{t\to\infty}\mathcal{W}_{1}(P^{*}_{s}\mu,P^{*}_{s}P^{*}_{t}\delta_{0})\leq\lim_{t\to\infty}Ce^{-\gamma s}\mathcal{W}_{1}(\mu,P_{t}^{*}\delta_{0})=0,\ \ s\geq 0.$
This, in turn, gives us
$\mathcal{W}_{1}(P_{s}^{*}\mu,\mu)$ $\leq\lim_{t\to\infty}(\mathcal{W}_{1}(P^{*}_{s}\mu,P^{*}_{s}P^{*}_{t}\delta_{0})+\mathcal{W}_{1}(P^{*}_{s}P^{*}_{t}\delta_{0},P^{*}_{t}\delta_{0})+\mathcal{W}_{1}(P^{*}_{t}\delta_{0},\mu))$
$\leq\lim_{t\to\infty}\mathcal{W}_{1}(P^{*}_{s}P^{*}_{t}\delta_{0},P^{*}_{t}\delta_{0})=\lim_{t\to\infty}\mathcal{W}_{1}(P^{*}_{t+s}\delta_{0},P^{*}_{t}\delta_{0})=0.$
As a result, $\mu$ is an invariant probability measure. Furthermore, by Theorem 3.3, we have
$\mathcal{W}_{1}(P^{*}_{t}\nu_{0},\mu)=\mathcal{W}_{1}(P^{*}_{t}\nu_{0},P^{*}_{t}\mu)\leq Ce^{-\gamma t}\mathcal{W}_{1}(\nu_{0},\mu),\ t\geq 0;$
this establishes (4.7) and completes the proof. ∎
To proceed, we use Theorem 3.6 to derive exponential ergodicity for the semigroup $P_{\cdot}^{*}$ of (1.1).
Theorem 4.5.
Given the assumptions of Theorem 3.6, $P^{*}_{t}$ admits an invariant measure $\mu\in\mathcal{P}_{1}(\mathbb{R}^{d})$ such that
$\mathcal{W}_{1}(P_{t}^{*}\nu_{0},\mu)\leq K\mathcal{W}_{\rho_{1}}(\mu,\nu_{0})e^{-ct},\ \ \forall\nu_{0}\in\mathcal{P}_{1}(\mathbb{R}^{d}),$ (4.10)
where $K$, $c$, and $\rho_{1}$ are defined in Theorem 3.6.
Proof.
We first note that Lemma 3.8 says that $\sup_{t\geq 0}\mathbb{E}[V(X_{t}(0))]<\infty$, where $X_{t}(0)$ denotes the solution to (1.1) with initial condition $X_{0}=0$. This, together with the fact that $f(x)\leq f(R_{2})$ with the function $f$ defined in (3.30) implies that
$\sup_{t\geq 0}\mathbb{E}[\rho_{1}(X_{t}(0),0)]\leq\sup_{t\geq 0}\mathbb{E}[f(|X_{t}(0)|)(1+\varepsilon V(X_{t}(0))+\varepsilon V(0))]<\infty.$
Next, we use (4.8) and (3.26) to obtain
$\lim_{t\to\infty}\sup_{s\geq 0}\mathcal{W}_{1}(P_{t+s}^{*}\delta_{0},P_{t}^{*}\delta_{0})$ $=\lim_{t\to\infty}\sup_{s\geq 0}\mathcal{W}_{1}(\mathcal{L}(X_{t}(P_{s}\delta_{0})),\mathcal{L}(X_{t}(0)))$ (4.11)
$\leq\lim_{t\to\infty}\sup_{s\geq 0}Ke^{-ct}\mathcal{W}_{\rho_{1}}(P_{s}^{*}\delta_{0},\delta_{0})$
$=\lim_{t\to\infty}\sup_{s\geq 0}K\mathbb{E}[\rho_{1}(X_{s}(0),0)]e^{-ct}$
$=0.$
The rest of the argument is similar to that in proof of Theorem 4.4. ∎
Appendix A Proofs of Several Lemmas
Proof of Lemma 2.1.
We use the idea in the proof of (Zimmer, , 2017, Lemma 2). First we consider the map $Z_{t}\to|Z_{t}|^{2}=r_{t}^{2}$. Then by Itô’s formula, we have
$\mathrm{d}(|Z_{t}|^{2})=$ $(2\langle Z_{t},\beta_{t}\rangle+\mathrm{rc}^{2}(U_{t})\mathrm{tr}(\alpha_{t}\alpha_{t}^{\top})+\mathrm{sc}^{2}(U_{t})\mathrm{tr}(\Delta_{t}\Delta_{t}^{\top}))\mathrm{d}t$ (A.1)
$+2\mathrm{sc}(U_{t})\langle Z_{t},\Delta_{t}\mathrm{d}B_{t}^{1}\rangle+2\mathrm{rc}(U_{t})\langle Z_{t},\alpha_{t}\mathrm{d}B_{t}^{2}\rangle$
First note that by (Wang, , 2018, Theorem 2.1), there exists a function $F(t)$ such that $\sup_{s\in[0,T]}\mathbb{E}[|Z_{s}|]\leq F(T)$ for all $T>0$. Now for any $\varepsilon\in(0,1)$, we define the $C^{2}$ function $g_{\varepsilon}(r):[0,\infty)\to(0,\infty)$ by
$g_{\varepsilon}(r)=\begin{cases}-\frac{r^{2}}{8\varepsilon^{3/2}}+\frac{3r}{4\varepsilon^{1/2}}+\frac{3\varepsilon^{1/2}}{8},&r<\varepsilon\\ \sqrt{r},&r\geq\varepsilon.\end{cases}$ (A.2)
Then we have
$\frac{dg_{\varepsilon}(r)}{dr}=\begin{cases}-\frac{r}{4\varepsilon^{3/2}}+\frac{3}{4\varepsilon^{1/2}},&r<\varepsilon,\\ \frac{1}{2\sqrt{r}},&r\geq\varepsilon,\end{cases}\quad\text{ and }\quad\frac{d^{2}g_{\varepsilon}(r)}{dr^{2}}=\begin{cases}-\frac{1}{4\varepsilon^{3/2}},&r<\varepsilon,\\ -\frac{1}{4r^{3/2}},&r\geq\varepsilon.\end{cases}$
Clearly $g_{\varepsilon}(r)\in C^{2}(\mathbb{R_{+}})$. Now fix $\delta>0$ and let $\varepsilon^{2}<\delta/2$, then we have by (A.1) and Itô’s formula
$g_{\varepsilon}$ $(|Z_{t}|^{2})-g_{\varepsilon}(|Z_{0}|^{2})$
$=$ $\,\!\int_{0}^{t}I_{\{|Z_{s}|^{2}\geq\varepsilon\}}\!\bigg(\langle e_{s},\beta_{s}\rangle+\frac{\mathrm{sc}^{2}(U_{s})}{2r_{s}}(\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2})+\frac{\mathrm{rc}^{2}(U_{s})}{2r_{s}}(\mathrm{tr}(\alpha_{s}\alpha_{s}^{\top})-|\alpha_{s}^{\top}e_{s}|^{2})\bigg)\mathrm{d}s$
$+\int_{0}^{t}I_{\{|Z_{s}|^{2}\geq\varepsilon\}}\mathrm{sc}(U_{s})\langle e_{s},\Delta_{s}\mathrm{d}B_{s}^{1}\rangle+\int_{0}^{t}\mathrm{rc}(U_{s})\langle e_{s},\alpha_{s}\mathrm{d}B_{s}^{2}\rangle$
$+\int_{0}^{t}I_{\{|Z_{s}|^{2}<\varepsilon\}}\left(\left(-\frac{r_{s}^{2}}{4\varepsilon^{3/2}}+\frac{3}{4\varepsilon^{1/2}}\right)\left(\langle Z_{s},\beta_{s}\rangle+\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})\right)-\frac{1}{2\varepsilon^{3/2}}|\Delta_{s}^{\top}Z_{s}|^{2}\right)\mathrm{d}s$
$+\int_{0}^{t}I_{\{|Z_{s}|^{2}<\varepsilon\}}\left(-\frac{r_{s}^{2}}{4\varepsilon^{3/2}}+\frac{3}{4\varepsilon^{1/2}}\right)\langle Z_{s},\Delta_{s}\mathrm{d}B_{s}^{1}\rangle.$ (A.3)
Note that in the above, we used the facts that $\mathrm{rc}(U_{t})=0$ and $\mathrm{sc}(U_{t})=1$ when $r_{t}<\delta/2$.
We will derive (2.1) by passing to the limit when $\varepsilon\to 0$ in (A). To this end, note that
$\frac{1}{2\varepsilon^{1/2}}\leq-\frac{r}{4\varepsilon^{3/2}}+\frac{3}{4\varepsilon^{1/2}}\leq\frac{3}{4\varepsilon^{1/2}}$ (A.4)
for $0\leq r\leq\varepsilon$. Using Assumption 1.1 along with the facts that the function $\kappa$ is bounded and that $\mathcal{W}_{1}(\mu_{t},\nu_{t})\leq\mathbb{E}[|Z_{t}|]\leq F(T)$ for any $t\in[0,T]$, we get
$I_{\{|Z_{t}|^{2}<\varepsilon\}}|\langle Z_{t},\beta_{t}\rangle|\leq I_{\{|Z_{t}|^{2}<\varepsilon\}}(\kappa(|Z_{t}|)|Z_{t}|^{2}+L_{1}\mathcal{W}_{1}(\mu_{t},\nu_{t})|Z_{t}|)\leq\kappa_{0}\varepsilon+L_{1}F(T)\varepsilon^{1/2}.$ (A.5)
Furthermore, Assumption 1.2 along with the fact that $\|A\|^{2}\leq\mathrm{tr}(AA^{\top})\leq d\|A\|^{2}$ for any $A\in\mathbb{R}^{d\times d}$ shows that on the set $\{|Z_{t}|^{2}<\varepsilon\}$
$\mathrm{tr}(\Delta_{t}\Delta_{t}^{\top})\leq d\|\Delta_{t}\|^{2}\leq dL_{2}|Z_{t}|^{2}\leq dL_{2}\varepsilon.$ (A.6)
Similarly, Assumptions 1.1 and 1.2 imply that on the set $\{|Z_{t}|^{2}<\varepsilon\}$, we have
$|\Delta_{t}^{\top}Z_{t}|^{2}\leq|Z_{t}|^{2}\|\Delta_{t}\|^{2}\leq L_{2}\varepsilon^{2}.$ (A.7)
Combining equations (A.4)–(A.7), we derive
$I_{\{|Z_{s}|^{2}<\varepsilon\}}\left|\bigg(-\frac{r_{s}^{2}}{4\varepsilon^{3/2}}+\frac{3}{4\varepsilon^{1/2}}\bigg)\left(\langle Z_{s},\beta_{s}\rangle+\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})\right)-\frac{1}{4\varepsilon^{3/2}}|\Delta_{s}^{\top}Z_{s}|^{2}\right|\leq K\varepsilon^{1/2}\leq K.$
for some positive constant $K$ independent of $\varepsilon$ and $s\in[0,t]$. As a result, it follows from the dominated convergence theorem that as $\varepsilon\to 0$
$\int_{0}^{t}I_{\{|Z_{s}|^{2}<\varepsilon\}}\bigg(\bigg(-\frac{r_{s}^{2}}{4\varepsilon^{3/2}}+\frac{3}{4\varepsilon^{1/2}}\bigg)\left(\langle Z_{s},\beta_{s}\rangle+\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})\right)-\frac{1}{4\varepsilon^{3/2}}|\Delta_{s}^{\top}Z_{s}|^{2}\bigg)\mathrm{d}s\to 0,\ \ \ \text{a.s.}$ (A.8)
Equations (A.4) and (A.7) also enable us to derive
$\mathbb{E}\bigg[\int_{0}^{t}I_{\{|Z_{s}|^{2}<\varepsilon\}}\left(-\frac{r_{s}^{2}}{4\varepsilon^{3/2}}+\frac{3}{4\varepsilon^{1/2}}\right)^{2}|\Delta_{s}^{\top}Z_{s}|^{2}\mathrm{d}s\bigg]\leq Kt\varepsilon\to 0$
as $\varepsilon\to 0$. Therefore there exists a subsequence $\{\varepsilon_{k}\}$ where $\varepsilon_{k}\to 0$ as $k\to\infty$ such that
[…]
as $\varepsilon_{k}\to 0$.
Using similar arguments as those for (A.8) and (A.9), we can show that along a subsequence of $\{\varepsilon_{k}\}$ (still denoted by $\{\varepsilon_{k}\}$), we have
$\!\int_{0}^{t}I_{\{|Z_{s}|^{2}\geq\varepsilon\}}\!\bigg(\langle e_{s},\beta_{s}\rangle+\frac{\mathrm{sc}^{2}(U_{s})}{2r_{s}}(\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2})+\frac{\mathrm{rc}^{2}(U_{s})}{2r_{s}}(\mathrm{tr}(\alpha_{s}\alpha_{s}^{\top})-|\alpha_{s}^{\top}e_{s}|^{2})\bigg)\mathrm{d}s$
$\hskip 124.30447pt+\int_{0}^{t}I_{\{|Z_{s}|^{2}\geq\varepsilon\}}\mathrm{sc}(U_{s})\langle e_{s},\Delta_{s}\mathrm{d}B_{s}^{1}\rangle+\int_{0}^{t}I_{\{|Z_{s}|^{2}\geq\varepsilon\}}\mathrm{rc}(U_{s})\langle e_{s},\alpha_{s}\mathrm{d}B_{s}^{2}\rangle$
$\to\bigg[\!\int_{0}^{t}I_{\{|Z_{s}|\neq 0\}}\!\bigg(\langle e_{s},\beta_{s}\rangle+\frac{\mathrm{sc}^{2}(U_{s})}{2r_{s}}(\mathrm{tr}(\Delta_{s}\Delta_{s}^{\top})-|\Delta_{s}^{\top}e_{s}|^{2})+\frac{\mathrm{rc}^{2}(U_{s})}{2r_{s}}(\mathrm{tr}(\alpha_{s}\alpha_{s}^{\top})-|\alpha_{s}^{\top}e_{s}|^{2})\bigg)\mathrm{d}s$
$\hskip 124.30447pt+\int_{0}^{t}I_{\{|Z_{t}|\neq 0\}}\mathrm{sc}(U_{s})\langle e_{s},\Delta_{s}\mathrm{d}B_{s}^{1}\rangle+\int_{0}^{t}I_{\{|Z_{t}|\neq 0\}}\mathrm{rc}(U_{s})\langle e_{s},\alpha_{s}\mathrm{d}B_{s}^{2}\rangle$
a.s. as $\varepsilon_{k}\to 0$. This, together with (A.8) and (A.9), gives (2.1). ∎
Proof of Lemma 2.2.
For simplicity of notation, we denote $\sigma=\sigma(y)$, $u=u(x,y),\Delta=\Delta(x,y)$, and $\alpha=\alpha(x,y)$ in this proof. Noting $\alpha=\Delta+2\sigma uu^{\top}$, we have
$\mathrm{tr}(\alpha\alpha^{\top})-|\alpha^{\top}e|^{2}=$ $\mathrm{tr}(\Delta\Delta^{\top})+4\mathrm{tr}(\Delta(\sigma uu^{\top})^{\top})+\frac{4}{|\sigma^{-1}z|^{2}}\mathrm{tr}(zz^{\top})-|\Delta^{\top}e+2uu^{\top}\sigma^{\top}e|^{2}$
$=$ $\mathrm{tr}(\Delta\Delta^{\top})+\frac{4}{|\sigma^{-1}z|^{2}}\mathrm{tr}(\Delta(zz^{\top}(\sigma^{-1})^{\top})^{\top})+\frac{4}{|\sigma^{-1}z|^{2}}|z|^{2}$
$-|\Delta^{\top}e|^{2}-4\mathrm{tr}(\Delta e(uu^{\top}\sigma^{\top}e)^{\top})-4|uu^{\top}\sigma^{\top}e|^{2}$
$=$ $\mathrm{tr}(\Delta\Delta^{\top})+\frac{4}{|\sigma^{-1}z|^{2}}\mathrm{tr}(\Delta^{\top}(zz^{\top}(\sigma^{-1})^{\top}))+\frac{4}{|\sigma^{-1}z|^{2}}|z|^{2}$
$-|\Delta^{\top}e|^{2}-4\mathrm{tr}(\Delta^{\top}zz^{\top}(\sigma^{-1})^{\top})-\frac{4}{|\sigma^{-1}z|^{2}}4|z|^{2}$
$=$ $\mathrm{tr}(\Delta\Delta^{\top})-|\Delta^{\top}e|^{2},$
where we used the definitions $u=\sigma^{-1}z/|\sigma^{-1}z|$, and $e=z/|z|$.
Assumption 1.2 gives us
$\big|\mathrm{tr}(\Delta\Delta^{\top})-|\Delta^{\top}e|^{2}|\leq|\Delta|^{2}+|\Delta^{\top}e\big|^{2}\leq 2|\Delta|^{2}\leq 2d\|\Delta\|^{2}\leq 2dL_{2}|x-y|^{2}.$
The proof is complete. ∎
Proof of Lemma 4.1.
First note that Assumption 1.3 gives us $|b(0,\delta_{0})|\leq L_{3}<\infty$, where $\delta_{0}$ is the Dirac measure at 0. Denote $X_{t}=X_{t}(0)$ throughout the proof. Then Itô’s formula gives
$\mathrm{d}(|X_{t}|^{2})=\left(2\langle X_{t},b(X_{t},\mu_{t})\rangle+\mathrm{tr}(\sigma(X_{t})\sigma(X_{t})^{\top})\right)\mathrm{d}t+2\langle X_{t},\sigma(X_{t})\mathrm{d}W_{t}\rangle.$
Consider the function $g\in C^{2}([0,\infty))$
$g(s):=\begin{cases}\sqrt{s},&\text{ if }s\geq 1,\\ -\frac{1}{8}s^{2}+\frac{3}{4}s+\frac{3}{8},&\text{ if }0\leq s<1,\end{cases}$
which has derivatives
[…]
Put $\tilde{e}_{t}=\frac{X_{t}}{|X_{t}|}$ for $X_{t}\neq 0$. Then we have by Itô’s formula that
$\mathrm{d}$ $(g(|X_{t}|^{2}))$
$=I_{\{|X_{t}|<1\}}\left(\left(-\frac{|X_{t}|^{2}}{4}+\frac{3}{4}\right)\left(2\langle X_{t},b(X_{t},\mu_{t})\rangle+\mathrm{tr}(\sigma(X_{t})\right)\sigma(X_{t})^{\top})-\frac{1}{2}|\sigma(X_{t})^{\top}X_{t}|^{2}\right)\mathrm{d}t$
$\ \ +I_{\{|X_{t}|\geq 1\}}\left(\langle\tilde{e}_{t},b(X_{t},\mu_{t})\rangle+\frac{1}{2|X_{t}|}\mathrm{tr}(\sigma(X_{t})\sigma(X_{t})^{\top})-\frac{1}{2}|\sigma(X_{t})^{\top}\tilde{e}_{t}|^{2}\right)\mathrm{d}t+\mathrm{d}M_{t},$
where $M_{t}$ is a martingale. Then, it follows that
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[g(|X_{t}|^{2})]$
$=\mathbb{E}\left[I_{\{|X_{t}|<1\}}\left(\left(-\frac{|X_{t}|^{2}}{4}+\frac{3}{4}\right)\left(2\langle X_{t},b(X_{t},\mu_{t})\rangle+\mathrm{tr}(\sigma(X_{t})\right)\sigma(X_{t})^{\top})-\frac{1}{2}|\sigma(X_{t})X_{t}|^{2}\right)\right]$
$\ \ +\mathbb{E}\left[I_{\{|X_{t}|\geq 1\}}\left(\langle\tilde{e}_{t},b(X_{t},\mu_{t})\rangle+\frac{1}{2|X_{t}|}(\mathrm{tr}(\sigma(X_{t})\sigma(X_{t})^{\top})-\frac{1}{2}|\sigma(X_{t})^{\top}\tilde{e}_{t}|^{2})\right)\right]$
$\leq\mathbb{E}\left[I_{\{|X_{t}|\geq 1\}}\langle\tilde{e}_{t},b(X_{t},\mu_{t})\rangle+I_{\{|X_{t}|<1\}}\frac{3-|X_{t}|^{2}}{2}\langle X_{t},b(X_{t},\mu_{t})\rangle+K_{0}\right],$ (A.10)
where $K_{0}=\sup\{\mathrm{tr}(\sigma(x)\sigma(x)^{\top})\}$, which is finite by Assumption 1.4. Now by Assumption 1.1, we have
$\langle X_{t},b(X_{t},\mu_{t})\rangle$ $=\langle X_{t},b(0,\delta_{0})\rangle+\langle X_{t},b(X_{t},\mu_{t})-b(0,\delta_{0})\rangle$
$\leq\kappa(|X_{t}|)|X_{t}|^{2}+L_{1}\mathcal{W}_{1}(\mu_{t},\delta_{0})|X_{t}|+\langle X_{t},b(0,\delta_{0})\rangle$
$\leq\kappa(|X_{t}|)|X_{t}|^{2}+L_{1}\mathbb{E}[|X_{t}|]|X_{t}|+|X_{t}||b(0,\delta_{0})|.$
Thus
$I_{\{|X_{t}|\geq 1\}}\langle\tilde{e}_{t},b(X_{t},\mu_{t})\rangle\leq I_{\{|X_{t}|\geq 1}\kappa(|X_{t}|))|X_{t}|+L_{1}\mathbb{E}[|X_{t}|]+|b(0,\delta_{0})|,$
and
$I_{\{|X_{t}|<1\}}\frac{3-|X_{t}|^{2}}{2}\langle X_{t},b(X_{t},\mu_{t})\rangle$ $\leq I_{\{|X_{t}|<1\}}\frac{3|X_{t}|-|X_{t}|^{3}}{2}(\kappa(|X_{t}|)|X_{t}|+L_{1}\mathbb{E}[|X_{t}|]+|b(0,\delta_{0})|)$
$\leq I_{\{|X_{t}|<1\}}(\kappa_{1}+|b(0,\delta_{0})|+L_{1}\mathbb{E}[|X_{t}|]),$
where $\kappa_{1}=\sup_{x\in[0,1]}|\kappa(x)|$; note also that we used the fact that $3x-x^{3}\leq 2$ for $x\in[0,1]$ in the last inequality. Now plugging these observations into (A) leads to
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[g(|X_{t}|^{2})]$ $\leq\mathbb{E}\left[I_{\{|X_{t}|\geq 1\}}\left(\kappa(|X_{t}|)|X_{t}|+L_{1}\mathbb{E}[|X_{t}|]+|b(0,\delta_{0})|\right)\right]$
$\quad+\mathbb{E}\left[I_{\{|X_{t}|\leq 1\}}\left(\kappa_{1}+|b(0,\delta_{0})|+L_{1}\mathbb{E}[|X_{t}|]\right)\right]+K_{0}$
$\leq\mathbb{E}\left[I_{\{|X_{t}|\geq 1\}}\kappa(|X_{t}|)|X_{t}|+L_{1}\mathbb{E}[|X_{t}|]+C_{1}\right],$
where $C_{1}=K_{0}+\kappa_{1}+|b(0,\delta_{0})|$. Now using $R_{0}$ from (4.1) we can define $\kappa_{R_{0}}=\sup_{x\in[0,R_{0}]}|\kappa(x)|$ and $C_{2}=C_{1}+\kappa_{R_{0}}R_{0}$. Then
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[g(|X_{t}|^{2})]$ $\leq\mathbb{E}\left[I_{\{|X_{t}|>R_{0}\}}(-L_{1}-K)|X_{t}|+\kappa_{R_{0}}R_{0}+L_{1}\mathbb{E}[|X_{t}|]+C_{1}\right]$
$\leq\mathbb{E}\left[I_{\{|X_{t}|>R_{0}\}}(-L_{1}-K)|X_{t}|+L_{1}\mathbb{E}[|X_{t}|]+C_{2}\right]$
$=\mathbb{E}\left[(-L_{1}-K)|X_{t}|+I_{\{|X_{t}|\leq R_{0}\}}(L_{1}+K)|X_{t}|+L_{1}\mathbb{E}[|X_{t}|]+C_{2}\right]$
$\leq\mathbb{E}\left[(-L_{1}-K)|X_{t}|+I_{\{|X_{t}|\leq R_{0}\}}(L_{1}+K)R_{0}+L_{1}\mathbb{E}[|X_{t}|]+C_{2}\right]$
$\leq-K\mathbb{E}[|X_{t}|]+C_{3},$
where $C_{3}=C_{2}+(L_{1}+K_{1})R_{0}$. Furthermore, using the elementary inequality $x\geq-\frac{1}{8}x^{4}+\frac{3}{4}x^{2}$ for $x\in[0,1]$, we have
$|X_{t}|$ $=I_{\{|X_{t}|<1\}}|X_{t}|+I_{\{|X_{t}|\geq 1\}}|X_{t}|$
$\geq I_{\{|X_{t}|<1}\left(-\frac{1}{8}|X_{t}|^{4}+\frac{3}{4}|X_{t}|^{2}\right)+I_{\{|X_{t}|\geq 1\}}\sqrt{|X_{t}|^{2}}$
$\geq g(|X_{t}|^{2})-\frac{3}{8}.$
Consequently, it follows that
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[g(|X_{t}|^{2})]\leq-K\mathbb{E}[g(|X_{t}|^{2})+C_{4}$
where $C_{4}=C_{3}+\frac{3}{8}K$. By Gronwall’s inequality (Chen and Li, , 1989, lemma 2.2) we have
$\mathbb{E}[g(|X_{t}|^{2})]\leq\frac{C_{4}}{K}(1-e^{-Kt})<\infty,\ \ \text{ for }t\geq 0.$
Last, we note that $x\leq g(x^{2})+1$ for all $x\in\mathbb{R}$ and hence
$\sup_{t\geq 0}\mathbb{E}[|X_{t}|]\leq 1+\sup_{t\geq 0}\mathbb{E}[g(|X_{t}|^{2})]<\infty.$
The proof is complete. ∎